### v1

# Brauer groups of abelian varieties over fields of finite characteristic

Abstract
We study the Brauer group of an abelian variety $A$ over an algebraically closed field of characteristic $p$ focusing on the $p$-primary torsion, the key part of which is a certain quasi-algebraic unipotent group ${\bf U}_{A}$. We determine its dimension and obtain a sharp upper bound for its $p$-exponent. The isogeny class of ${\bf U}_{A}$ is classified for abelian varieties $A$ of dimension at most 3. For principally polarised abelian varieties we compute the dimension of the $p$-torsion subgroup of ${\bf U}_{A}$ in terms of the Ekedahl–Oort type of $A$.
Introduction
Let $k$ be an algebraically closed field of characteristic $p\geqslant 0$. In this paper we are interested in the following question: what is the Brauer group of an abelian variety $A$ over $k$? In particular, what is the $p$-primary torsion subgroup ${\rm{Br}}(A)[p^{\infty}]$ when $p>0$?
Let us start by recalling the structure of the Brauer group of an arbitrary smooth proper variety $X$ over $k$. In [GB, III] Grothendieck proved that ${\rm{Br}}(X)$ is a torsion group and computed ${\rm{Br}}(X)[\ell^{\infty}]$ for primes $\ell\neq p$. The exactness of the Kummer sequence with coefficients $\mu_{\ell^{n}}$ for the étale topology gives an isomorphism
${\rm{Br}}(X)[\ell^{\infty}]\cong({\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})^{b_{2}-\rho}\oplus{\rm H}^{3}_{\rm{\acute{e}t}}(X,{\mathbb{Z}}_{\ell}(1))[\ell^{\infty}],$
where $b_{2}={\rm{dim}}_{{\mathbb{Q}}_{\ell}}{\rm H}^{2}_{\rm{\acute{e}t}}(X,{\mathbb{Q}}_{\ell})$ is the second $\ell$-adic Betti number, $\rho={\rm{rk}}\,{\rm NS\,}(X)$ is the Picard number, and ${\rm H}^{3}_{\rm{\acute{e}t}}(X,{\mathbb{Z}}_{\ell}(1))[\ell^{\infty}]$ is finite, see [CTS21, Prop. 5.2.9]. When $X=A$ is an abelian variety of dimension $g>0$, we have $b_{2}=g(2g-1)$ and ${\rm H}^{3}_{\rm{\acute{e}t}}(A,{\mathbb{Z}}_{\ell}(1))[\ell^{\infty}]=0$. In particular, the abelian group ${\rm{Br}}(A)[\ell^{\infty}]$ is an invariant of the isogeny class of $A$.
[…]
${\rm{Br}}(X)[p^{\infty}]\cong({\mathbb{Q}}_{p}/{\mathbb{Z}}_{p})^{r_{2}-\rho}\oplus{\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}],$ (1)
see [P24, Theorem A.1]. In contrast with the $\ell$-adic case, the group ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$ can be infinite. A phenomenon specific to finite characteristic is the existence of a natural commutative quasi-algebraic group over $k$ whose group of $k$-points is ${\rm H}^{3}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$. This goes back to Milne [Mil76, Cor. 2.7 (a)], see [Ber81, Cor. 2.10], [IR83, Lemme IV.3.2.2], [Mil86, Lemma 1.8]. Quasi-algebraic groups were introduced by Serre [Ser60, §1.2] as equivalence classes of algebraic groups with a common set of $k$-points, where two groups $G$ and $H$ are called equivalent if there is a morphism $G\to H^{(p^{n})}$ which is a bijection on $k$-points.
When $X=A$ is an abelian variety, the attached quasi-algebraic group is connected, see [YY, Thm. 1.5] based on [IR83, Thm. II.3.4] (“survie du cœur”). Hence this is a commutative connected unipotent quasi-algebraic group, which we denote by ${\bf U}_{A}$. In particular, the group of $k$-points ${\bf U}_{A}(k)\cong{\rm H}^{3}(A,{\mathbb{Z}}_{p}(1))[p^{\infty}]$, the dimension ${\rm{dim}}({\bf U}_{A})$, and the $p$-exponent $\operatorname{\mathnormal{p}-exp}({\bf U}_{A})$ (the least integer $m$ such that $p^{m}{\bf U}_{A}=0$) are all well-defined. By general theory [Ser59, VII.10], ${\bf U}_{A}$ is isogenous to a product of truncated Witt (quasi-algebraic) groups $\prod_{i=1}^{r}{\bf W}_{n_{i}}$, for some positive integers $n_{1}\leqslant\ldots\leqslant n_{r}$. We have
${\rm{dim}}({\bf U}_{A})=\sum_{i=1}^{r}n_{i},\quad{\rm{dim}}({\bf U}_{A}[p])=r,\quad\operatorname{\mathnormal{p}-exp}({\bf U}_{A})=n_{r}.$
Note that if $\operatorname{\mathnormal{p}-exp}({\bf U}_{A})=1$, then ${\bf U}_{A}\simeq{\bf G}_{a,k}^{d}$, see [Ser59, VII.2, Prop. 11].
The guiding question of this paper is: what sequences $(n_{1},\ldots,n_{r})$ occur for abelian varieties of dimension $g$ over an algebraically closed field $k$ of positive characteristic $p$?
Using results of Crew and deep results of Ekedahl, together with the fact that the groups ${\rm H}^{i}_{\rm cris}(A/W)$ are torsion-free and the Hodge-to-de Rham spectral sequence for $A$ degenerates at the first page, one obtains [YY, Theorem 1.1, 1.6]
${\rm{dim}}({\bf U}_{A})=\frac{g(g-1)}{2}-\sum_{0\leqslant\lambda<1}(1-\lambda)m_{\lambda},$ (2)
where $g={\rm{dim}}(A)$ and $m_{\lambda}$ is the multiplicity of slope $\lambda$ in ${\rm H}^{2}_{\rm cris}(A/W)$. In particular, ${\rm{dim}}({\bf U}_{A})$ depends only on the slopes of $A$ and thus is an invariant of the isogeny class of $A$. The meaning of the terms of (2) is the following: ${\rm{dim}}_{k}{\rm H}^{0}(A,\Omega^{2}_{A/k})=g(g-1)/2$ is the dimension of the tangent space of the formal Brauer group $\widehat{{\rm{Br}}}(A)$, the sum is the dimension of the maximal quotient of $\widehat{{\rm{Br}}}(A)$ of finite height, and ${\rm{dim}}({\bf U}_{A})$ is the dimension of the unipotent part of $\widehat{{\rm{Br}}}(A)$. Thus (2) implies the smoothness of the formal Brauer group of any abelian variety over $k$. Note that smoothness of $\widehat{{\rm{Br}}}(A)$ is a particular case of a general criterion of smoothness of Artin–Mazur formal groups obtained by the first named author in [Gr25b].
From (2) we see that ${\rm{dim}}({\bf U}_{A})\leqslant g(g-1)/2$ with equality if and only if $g=1$ or $A$ is supersingular (Remark 3.2). For $g=2$, it gives ${\bf U}_{A}=0$ unless $A$ is supersingular, in which case ${\bf U}_{A}\cong{\bf G}_{a,k}$, a fact that goes back to Illusie whose computation was based on the de Rham–Witt cohomology [Ill79, II.7]. In general, it is known that ${\bf U}_{A}=0$ when ${\rm H}^{2}(A,W{\mathcal{O}}_{A})$ is a finitely generated $W$-module, for example, when $A$ is ordinary (i.e., the $p$-rank is $g$) or almost ordinary (i.e., the $p$-rank is g1g-1), see [Ill83, Corollary 6.3.16], [YY, Theorem 3.7]. In fact, ${\bf U}_{A}=0$ if and only if $A$ is ordinary or almost ordinary, see Remark 3.2.
Let us now describe the main results of this paper and give its brief outline.
  • Developing the approach of [Gr25a] based on the Nygaard filtration, we express ${\bf U}_{A}$ in terms of the F-crystal ${\rm H}^{2}_{\rm cris}(A/W)$, see Corollary 1.7. This gives a quick proof that ${\rm{dim}}({\bf U}_{A})$ depends only on the isogeny class of $A$ (Proposition 2.6). From this we obtain a proof of formula (2) that uses only the classical Dieudonné theory and so avoids the deep theory of Ekedahl, see Theorem 3.1. At the end of §3.1 we give a different, short proof of (2) based on the smoothness of the formal Brauer group of $A$.
  • We interpret ${\rm{Br}}(A)$ in terms of homomorphisms of truncated Dieudonné modules. Namely, for $p\neq 2$ there is a natural isomorphism of ${\bf U}_{A}[p^{n}]$ and the quasi-algebraic group
    ${\rm{Hom}}_{\mathbb{E}}(M_{A}^{\vee}/p^{n},M_{A}/p^{n})^{\rm skew}/\left({\rm{Hom}}_{\mathbb{E}}(M_{A}^{\vee},M_{A})^{\rm skew}/p^{n}\right),$
    which is a subquotient of the quasi-algebraic group ${\rm{Hom}}_{W(k)}(M_{A}^{\vee}/p^{n},$ $M_{A}/p^{n})$, see Corollary 2.5. Here $M_{A}$ is the Dieudonné module of $A$, $\mathbb{E}$ is the Dieudonné ring (see §1.1) and $\rm skew$}’ stands for the subgroup of anti-self-dual elements. This gives a different way to reduce the determination of ${\bf U}_{A}$ to a computation with the Dieudonné module $M_{A}$.
  • We prove that the $p$-exponent of ${\bf U}_{A}$ is at most g1g-1, where $g={\rm{dim}}(A)$ and $p\neq 2$, see Theorem 3.7. We show that the bound is sharp by exhibiting, for every $g\geqslant 1$, a supergeneral (i.e. supersingular with $a$-number 1) abelian variety $A$ of dimension $g$ such that ${\bf U}_{A}$ is isogenous to $\prod_{n=1}^{g-1}{\bf W}_{n}$. Combining this with a result of D’Addezio we obtain that if $A$ is an abelian variety over a field $k_{0}$ finitely generated over ${\mathbb{F}}_{p}$, and $k=\overline{k_{0}}$, then the transcendental Brauer group ${\rm{Im}}[{\rm{Br}}(A)\to{\rm{Br}}(A_{k})]$ is a direct sum of a finite group and an abelian group annihilated by $p^{g-1}$ (Corollary 3.8). Moreover, if $A$ is supersingular, then the transcendental Brauer group of $A$ is annihilated by $p^{g-1}$, whereas if $A$ is ordinary or almost ordinary, then this group is finite.
  • We determine the isogeny class of ${\bf U}_{A}$ for arbitrary abelian threefolds (assuming $p\neq 2$): ${\bf U}_{A}\simeq{\bf G}_{a,k}^{n}$, where $0\leqslant n\leqslant 3$, unless $A$ is supergeneral, in which case $U_{A}$ is isogenous to ${\bf G}_{a,k}\times{\bf W}_{2}$, see Proposition 3.15. In particular, the isogeny class of ${\bf U}_{A}$ is not an invariant of the isogeny class of $A$.
  • We compute ${\rm{dim}}({\bf U}_{A}[p])$ for principally polarised abelian varieties $A$ of arbitrary dimension $g\geqslant 1$ in terms of the Ekedahl–Oort type of $A[p]$, see Theorem 4.1. From this we deduce that in the principally polarised case we have ${\rm{dim}}({\bf U}_{A})\leqslant r(r+1)/2$ and $\operatorname{\mathnormal{p}-exp}({\bf U}_{A})\leqslant r$, where $r={\rm{dim}}({\bf U}_{A}[p])$, see Corollary 4.7.
The work on this paper started when A.S. was a Shapiro Visitor in Pennsylvania State University. He is very grateful to Yuri Zarhin for many helpful discussions without which this paper would have never been written. L.G. received funding from the EU Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement 101126554. The work of Y.Y. was supported by the London School of Geometry and Number Theory through the Engineering and Physical Sciences Research Council grant EP/S021590/1. The authors are grateful to the Bernoulli Center at EPFL, where a large part of this work was carried out, for excellent working conditions and support. We would like to thank James Borger, George Boxer, Daniel Bragg, Vladimir Drinfeld, Mikhail Kapranov, and Alexander Petrov for their questions and helpful discussions.
1 Preliminaries
1.1 Notation
In this paper $k$ is an algebraically closed field of characteristic $p>0$. We write $W=W(k)$ for the ring of Witt vectors, and denote by $\sigma\colon W\to W$ the canonical lifting of the Frobenius $x\to x^{p}$ on $k$.
All group schemes in this paper are commutative. For a finite flat commutative group $k$-scheme $G$ we denote by $G^{\vee}$ its Cartier dual. If $A$ is an abelian variety, we write $A^{\vee}$ for the dual abelian variety.
The Dieudonné ring ${\mathbb{E}}=W_{\sigma}[F,V]$ is the ring of non-commutative polynomials in $F$ and $V$ with coefficients in $W$ subject to the relations $FV=VF=p$, $F\lambda=\lambda^{\sigma}F$, $V\lambda=\lambda^{\sigma^{-1}}V$, where $\lambda\in W$. An $\mathbb{E}$}-module $M$, which is free and finitely generated as a $W$-module, will be called a Dieudonné module. The dual Dieudonné module is defined as $M^{\vee}={\rm{Hom}}_{W}(M,W)$ with the action of $\mathbb{E}$} on $f\in M^{\vee}$ given by the rules $Ff(x)=\sigma f(Vx)$ and $Vf(x)=\sigma^{-1}f(Fx)$, for any $x\in M$.
1.2 Quasi-algebraic and pro-algebraic groups
Quasi-algebraic groups were introduced by Serre [Ser60] as a way to have an abelian category of group schemes where infinitesimal groups and radicial morphisms are ignored. This is useful whenever one is interested in working with $k$-points of algebraic groups, rather than algebraic groups themselves, as in our case. Serre’s construction can be interpreted as restricting a group scheme over $k$ to the category of perfect schemes over $k$.
Let $\pi\colon X\to{\rm{Spec}}(k)$ be a proper and smooth scheme over $k$. Let ${\rm{Spec}}(k)_{\rm perf}$ be the perfect site of $k$. The objects of ${\rm{Spec}}(k)_{\rm perf}$ are perfect $k$-schemes (that is, $k$-schemes such that the Frobenius map is an isomorphism) equipped with the étale topology. Let $(X/k)_{\rm perf}$ be the site such that the objects are pairs $(T,Y)$, where $T$ is a perfect $k$-scheme and $Y$ is étale over $X\times_{k}T$, equipped with the étale topology. By [Ber81, 2.7] we have a commutative diagram of topoi, where $X_{\rm Et}$ is the big étale site of $X$, and $\alpha$ and $\beta$ are natural restriction maps:
${X_{\rm fppf}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta_{X}}$$\scriptstyle{\pi_{\rm fppf}}$${X_{\rm Et}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{X}\ \ \ }$$\scriptstyle{\pi_{\rm Et}}$${(X/k)_{\rm perf}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${{\rm{Spec}}(k)_{\rm fppf}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta_{k}}$${{\rm{Spec}}(k)_{\rm Et}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{k}\ \ }$${{\rm{Spec}}(k)_{\rm perf}}$
[…]
Let ${\mathcal{G}}[p^{\infty}]$ be the category of perfect commutative affine $k$-groups annihilated by some power of $p$ that are perfections of group $k$-schemes of finite type. Let $\mathcal{P}$} be the category of pro-objects of ${\mathcal{G}}[p^{\infty}]$. The objects of ${\mathcal{G}}[p^{\infty}]$ (respectively, $\mathcal{P}$}) can be described as quasi-algebraic (respectively, pro-algebraic) $k$-groups in the sense of Serre [Ser60, §1, §2]. Serre shows that ${\mathcal{G}}[p^{\infty}]$ and $\mathcal{P}$} are abelian categories, where subobjects are closed subgroups [Ser60, §1.2, Prop. 5, §2.4, Prop. 7], see also [Ber81, 2.2 (a)]. An example of an object of ${\mathcal{G}}[p^{\infty}]$ is ${\bf W}_{n}=(\mathbb{W}_{n})^{\rm perf}$, where $\mathbb{W}_{n}$ is the group $k$-scheme of Witt vectors of length $n$. In particular, we have ${\bf G}_{a,k}:={\bf W}_{1}=({\mathbb{G}}_{a,k})^{\rm perf}$. An example of an object of $\mathcal{P}$} is ${\bf W}=\mathbb{W}^{\rm perf}$, see [Ser60, p. 13, Exemple 3].
Unipotent quasi-algebraic groups from F-crystals
Let $L$ be an F-crystal, i.e. a free finitely generated $W$-module with an injective $\sigma$-linear endomorphism $F$. Define
${{\rm F}^{1}_{N}{}}(L)=\{x\in L\text{ such that }p|F(x)\}.$
In particular, we have $pL\subset{{\rm F}^{1}_{N}{}}(L)\subset L$.
The F-crystal $L$ has a natural structure of a pro-algebraic group $\bf L$} such that ${\bf L}(k)=L$, so that $\bf L$} is isomorphic to ${\bf W}^{n}$ where $n={\rm{rk}}_{W}(L)$. Then $F$ naturally gives rise to a $k$-endomorphism ${\bf L}\to{\bf L}$ also denoted by $F$. It is the relative Frobenius $F_{{\bf L}/k}\colon{\bf L}\to{\bf L}$ followed by a linear map ${\bf L}\to{\bf L}$.
Thus ${{\rm F}^{1}_{N}{}}({\bf L})$ is a closed subgroup of $\bf L$}, and $F/p-1\colon{{\rm F}^{1}_{N}{}}({\bf L})\to{\bf L}$ is a morphism in the category of pro-algebraic $k$-groups $\mathcal{P}$}. But $\mathcal{P}$} is an abelian category, so
${\bf U}_{L}:={\rm{Coker}}[F/p-1\colon{{\rm F}^{1}_{N}{}}({\bf L})\to{\bf L}]$ (3)
is a pro-algebraic group.
Proposition 1.1
Let $L$ be an F-crystal. Then ${\bf U}_{L}$ is a connected unipotent quasi-algebraic group.
Proof. Let $n$ be the rank of the free finitely generated $W$-module $L$. By a standard argument [Ill79, Lemme II.5.6], the cokernel of $F/p-1\colon{{\rm F}^{1}_{N}{}}(L)\to L$ is a torsion abelian group. We claim that it is annihilated by $p^{r}$ for some $r\geqslant 1$. From this it follows that ${\bf U}_{L}$ is the quotient of the connected unipotent quasi-algebraic group $k$-scheme ${\bf W}_{r}^{n}$ by a closed subgroup. This implies the proposition.
It remains to prove the claim. We proceed as in [Gr25a, Lemma 5.1.3]. Since $L/{{\rm F}^{1}_{N}{}}(L)$ is killed by $p$, it is enough to prove that the cokernel of $F-p\colon L\to L$ has finite $p$-exponent. By the Manin–Dieudonné classification, it is enough to prove this for an F-crystal $M$ with a $W$-basis $e_{1},\ldots,e_{r}$ such that $Fe_{i}=e_{i+1}$ for $i=1,\ldots,r-1$ and $Fe_{r}=p^{s}e_{1}$, where $r>0$ and $s\geqslant 0$ are coprime integers. There are three possible cases:
(1) If $r=s=1$, then Fp=p(F/p1)F-p=p(F/p-1), and $(F/p-1)(\lambda e_{1})=(\sigma(\lambda)-\lambda)e_{1}$ is surjective by [Ill79, Lemme II.5.3].
(2) If $s>r$, then $p^{r}M\subset(F-p)M$. Indeed, if $a\in p^{r}M$, then the infinite sum $\sum_{i\geqslant 1}F^{i-1}p^{-i}a$ converges to some $x\in M$, and we have a=Fxpxa=Fx-px.
(3) If $s<r$, then $p^{s}M\subset(F-p)M$. Indeed, if $a\in p^{s}M$, then $\sum_{i\geqslant 0}F^{-i}p^{i}a$ converges to some $x\in M$, and we have a=Fxpxa=Fx-px. $\Box$
Definition 1.2
[…]
The $p$-exponent of a unipotent quasi-algebraic group $\bf U$} is the least non-negative integer $n$ such that $p^{n}{\bf U}=0$.
Corollary 1.3
Define ${\bf H}_{L,n}={\rm{Ker}}[F/p-1\colon{{\rm F}^{1}_{N}{}}({\bf L})/p^{n}\to{\bf L}/p^{n}]$. Then we have an exact sequence of quasi-algebraic $k$-groups
$0\to{\bf L}^{F=p}/p^{n}\to{\bf H}_{L,n}\to{\bf U}_{L}[p^{n}]\to 0.$ (4)
In particular, for $n$ greater or equal to the $p$-exponent of ${\bf U}_{L}$ the quasi-algebraic group ${\bf U}_{L}$ is naturally isomorphic to the cokernel of ${\bf L}^{F=p}/p^{n}\to{\bf H}_{L,n}$.
Proof. We have an exact sequence of pro-algebraic groups
$0\to{\bf L}^{F=p}\to{{\rm F}^{1}_{N}{}}({\bf L})\xrightarrow{F/p-1}{\bf L}\to{\bf U}_{L}\to 0.$
Breaking it into two short exact sequences and applying the snake lemma to the diagrams formed by multiplication by $p^{n}$, we obtain (4). $\Box$
Note that ${\bf L}^{F=p}$ is finitely generated as a ${\mathbb{Z}}_{p}$-module [Ill79, Lemme II.5.11], so ${\bf L}^{F=p}/p^{n}$ is finite. Thus the map ${\bf H}_{L,n}\to{\bf U}_{L}[p^{n}]$ is an isogeny of quasi-algebraic groups.
1.3 Quasi-algebraic groups from flat and crystalline cohomology
The sheaves $W_{n}\Omega^{i}$ on $X_{\rm{\acute{e}t}}$ can be lifted to sheaves on $(X/k)_{\rm perf}$, again denoted $W_{n}\Omega^{i}$, see [IR83, p. 192]. Formation of $W_{n}\Omega^{i}$ commutes with perfect base change [Ill79, Prop. I.1.9.2]. From this one deduces that the sheaves $R^{j}\pi_{*}(W_{n}\Omega^{i})$ on ${\rm{Spec}}(k)_{\rm perf}$ belong to ${\mathcal{G}}[p^{n}]$, that is, are represented by quasi-algebraic groups, see [IR83, Lemme IV.3.2.2]. Using the notation of [IR83] we denote these quasi-algebraic groups by ${\bf H}^{j}(X,W_{n}\Omega^{i})$. Considering the complex $W_{n}\Omega_{X/k}^{\bullet}$ as an object of the derived category of sheaves on $(X/k)_{{\rm perf}}$, we denote the quasi-algebraic group representing $R^{j}\pi_{*}(W_{n}\Omega_{X/k}^{\bullet})$ by ${\bf H}^{j}_{\rm cris}(X/W_{n})$. Passing to the limit we obtain pro-algebraic $k$-groups ${\bf H}^{j}(X,W\Omega^{i})$ and ${\bf H}^{j}_{\rm cris}(X/W)$. Since $k$ is algebraically closed, the group of $k$-points of ${\bf H}^{j}(X,W\Omega^{i})$ (respectively, ${\bf H}^{j}_{\rm cris}(X/W)$) is ${\rm H}^{j}(X,W\Omega^{i})$ (respectively, ${\rm H}^{j}_{\rm cris}(X/W)$), and the same holds with $W$ replaced by $W_{n}$. Moreover, for a perfect $k$-algebra $k^{\prime}$ we have
${\bf H}_{\rm cris}^{i}(X/W)(k^{\prime})=\Gamma({\rm{Spec}}(k^{\prime}),R^{j}\pi_{*}(W\Omega_{X/k}^{\bullet}))={\rm H}_{\rm cris}^{i}(X/W)\otimes_{W}W(k^{\prime}).$
When ${\rm H}_{\rm cris}^{i}(X/W)$ is torsion-free, this implies that ${\bf H}_{\rm cris}^{i}(X/W)$ is the pro-algebraic group ${\bf W}^{n}$ for some $n\geqslant 0$.
On all of the above pro-algebraic groups, the Frobenius $F$ acts as an endomorphism over $k$.
[…]
${\bf H}^{j}(X,{\mathbb{Z}}_{p}(1))\cong{\bf H}^{j-1}(X,W\Omega^{1}_{\rm log})\cong{\bf H}^{j-1}(X,W\Omega^{1})^{F=1}.$
As in [BMS, Def. 8.1] we define the first term ${{\rm F}^{1}_{N}{}}W\Omega^{\bullet}$ of the Nygaard filtration on the de Rham–Witt complex $W\Omega^{\bullet}$ as the complex
$0\to VWO_{X}\to W\Omega^{1}_{X}\to W\Omega^{2}_{X}\to\ldots$
Thus there is an exact sequence of complexes of pro-étale sheaves
$0\to{{\rm F}^{1}_{N}{}}W\Omega^{\bullet}_{X}\to W\Omega_{X}^{\bullet}\to O_{X}\to 0,$ (5)
hence a long exact sequence of pro-algebraic groups
$\ldots\to{\bf H}^{j-1}(X,O_{X})\to{{\rm F}^{1}_{N}{}}{\bf H}^{j}_{\rm cris}(X/W)\to{\bf H}^{j}_{\rm cris}(X/W)\to{\bf H}^{j}(X,O_{X})\to\ldots$ (6)
One can define maps $F/p:{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)\to{\bf H}_{\rm cris}^{i}(X/W)$ making the following diagram commute
${{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F/p}$$\scriptstyle{1}$${{\bf H}_{\rm cris}^{i}(X/W)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${{\bf H}_{\rm cris}^{i}(X/W)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}$${{\bf H}_{\rm cris}^{i}(X/W)}$
By [BMS, Prop. 8.4] the map F/p1F/p-1 fits into an exact sequence
$0\to W\Omega^{1}_{X,\log}[-1]\to{{\rm F}^{1}_{N}{}}W\Omega^{\bullet}_{X}\xrightarrow{F/p-1}W\Omega_{X}^{\bullet}\to 0$ (7)
This gives a long exact sequence of pro-algebraic groups
$\ldots\to{\bf H}^{j}(X,{\mathbb{Z}}_{p}(1))\to{{\rm F}^{1}_{N}{}}{\bf H}^{j}_{\rm cris}(X/W)\xrightarrow{F/p-1}{\bf H}^{j}_{\rm cris}(X/W)\to{\bf H}^{j+1}(X,{\mathbb{Z}}_{p}(1))\to\ldots$ (8)
see also [BL, Thm. 7.3.5] and [Gr25a, Thm. 4.5] for a different approach.
Definition 1.4
A smooth proper $k$-variety $X$ over an algebraically closed field $k$ of characteristic $p$ is called straight or Mazur–Ogus if ${\bf H}^{i}_{\rm cris}(X/W)$ is torsion-free for all $i\geqslant 0$, and the Hodge-to-de Rham spectral sequence of $X$ degenerates at the first page.
For example, abelian varieties and K3 surfaces are straight. For some properties of straight varieties, see [Ogus79, §1].
Proposition 1.5
A straight variety $X$ over an algebraically closed field $k$ of characteristic $p$ has the following properties.
(i) There is an exact sequence of pro-algebraic groups
$0\to{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)\to{\bf H}_{\rm cris}^{i}(X/W)\to{\bf H}^{i}(X,{\cal O}_{X})\to 0.$
Here ${{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)\subset{\bf H}_{\rm cris}^{i}(X/W)$ is the closed subgroup given by the condition that $p$ divides $F(x)$. The map
$F/p-1\colon{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)\to{\bf H}_{\rm cris}^{i}(X/W)$
is the obvious map.
(ii) There is an exact sequence of pro-algebraic groups
$0\to{\bf H}^{i}(X,{\mathbb{Z}}_{p}(1))/\mathrm{tors}\to{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)\xrightarrow{\frac{F}{p}-1}{\bf H}_{\rm cris}^{i}(X/W)\to{\bf H}^{i+1}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]\to 0.$
In particular, ${\bf H}^{i+1}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$ is the connected unipotent quasi-algebraic group attached to the $F$-crystal ${\bf H}_{\rm cris}^{i}(X/W)$.
(iii) There is a long exact sequence of quasi-algebraic groups
$\ldots\to{\bf H}^{i}(X,\mu_{p^{n}})\to{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)/p^{n}\xrightarrow{\frac{F}{p}-1}{\bf H}_{\rm cris}^{i}(X/W)/p^{n}\to{\bf H}^{i+1}(X,\mu_{p^{n}})\to\ldots$
Proof. (i) The map ${\bf H}_{\rm cris}^{i}(X/W)\to{\bf H}^{i}(X,{\cal O}_{X})$ is the composition of the natural maps
${\bf H}_{\rm cris}^{i}(X/W)\xrightarrow{\alpha}{\bf H}_{\rm dR}^{i}(X/W)\xrightarrow{\beta}{\bf H}^{i}(X,{\cal O}_{X}).$
Here $\alpha$ is surjective since ${\bf H}_{\rm cris}^{i+1}(X/W)$ is torsion-free, and $\beta$ is surjective by the degeneration of the Hodge-to-de Rham spectral sequence. Thus (6) gives the exact sequence in (i). The second property is a consequence of the Mazur–Ogus theorem, see also [Gr25a, Prop. 5.2.2].
(ii) The exact sequence follows from (8) since ${{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)$ and ${\bf H}_{\rm cris}^{i}(X/W)$ are both torsion-free, while the cokernel of F/p1F/p-1 is torsion, by a standard argument in the theory of F-isocrystals [Ill79, Lemme II.5.6]. Since ${\bf H}_{\rm cris}^{i}(X/W)$ is the natural pro-algebraic group of the F-crystal ${\rm H}_{\rm cris}^{i}(X/W)$, it remains to apply Proposition 1.1.
(iii) This is obtained by applying the derived functor $\otimes^{L}{\mathbb{Z}}/p^{n}$ to (7) and considering the associated long exact sequence, using that ${{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{i}(X/W)$ and ${\bf H}_{\rm cris}^{i}(X/W)$ are torsion-free. $\Box$
Remark 1.6
If $X$ is straight, then ${\rm H}^{0}(X,Z\Omega^{1}_{X})\to{\rm H}^{0}(X,\Omega^{1})$ is an isomorphism. (Here by definition $Z\Omega^{1}_{X}={\rm{Ker}}[d\colon\Omega^{1}_{X}\to\Omega^{2}_{X}]$.) Then [Ill79, Thm. II.5.14] gives ${\rm H}^{2}(X,{\mathbb{Z}}_{p}(1))\cong{\rm H}^{2}_{\rm cris}(X/W)^{F=p}$, so we have ${\rm H}^{2}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]=0$. From Proposition 1.5 (ii) we deduce the surjectivity of the map
$F/p-1:{{\rm F}^{1}_{N}{}}{}{\bf H}_{\rm cris}^{1}(X/W)\to{\bf H}_{\rm cris}^{1}(X/W).$
Thus the connected unipotent group attached to the F-crystal ${\bf H}_{\rm cris}^{1}(X/W)$ is zero. The exact sequence of Proposition 1.5 (iii) gives an isomorphism
${\bf H}^{2}(X,\mu_{p^{n}})\cong{\rm{Ker}}[{{\rm F}^{1}_{N}{}}{\bf H}_{\rm cris}^{2}(X/W)/p^{n}\xrightarrow{F/p-1}{\bf H}_{\rm cris}^{2}(X/W)/p^{n}].$ (9)
Corollary 1.7
Let $X$ be a straight variety over an algebraically closed field of characteristic $p>0$. Then the quotient of ${\rm{Br}}(X)$ by its maximal divisible subgroup is the group of $k$-points of the connected unipotent quasi-algebraic group associated to the F-crystal ${\rm H}^{2}_{\rm cris}(X/W)$.
In this paper we are primarily interested in the case when $X=A$ is an abelian variety.
Definition 1.8
Let $A$ be an abelian variety over an algebraically closed field $k$ of characteristic $p>0$. We denote by ${\bf U}_{A}$ the connected unipotent quasi-algebraic group associated to the F-crystal ${\rm H}^{2}_{\rm cris}(A/W)$.
In this case we have ${\rm H}^{2}_{\rm cris}(A/W)\cong\wedge^{2}_{W}{\rm H}^{1}_{\rm cris}(A/W)$, where ${\rm H}^{1}_{\rm cris}(A/W)$ is the Dieudonné module of the $p$-divisible group $A[p^{\infty}]$.
The first statement of the following corollary is [Gr25a, Cor. 5.2.4].
Corollary 1.9
Let $A$ be an abelian variety over an algebraically closed field $k$ of characteristic $p>0$, and let $i\geqslant 1$. The multiplication by $m$ map $m]:A\to A$ acts on ${\rm H}^{i}(A,{\mathbb{Z}}_{p}(1))/\mathrm{tors}$ as multiplication by $m^{i}$, on ${\rm H}^{i}(X,{\mathbb{Z}}_{p}(1))[p^{\infty}]$ as multiplication by $m^{i-1}$, and on ${\rm H}^{i}_{\rm fppf}(A,\mu_{p^{n}})$, for $i\leqslant 2$, and ${\rm{Br}}(A)$ as multiplication by $m^{i}$.
Proof. The statements concerning ${\rm H}^{i}(A,{\mathbb{Z}}_{p}(1))$ follow from the exact sequence of Proposition 1.5 (ii) and the well-known fact that $[m]$ acts as $m^{i}$ on ${\rm H}_{\rm cris}^{i}(A)\cong\wedge^{i}{\rm H}^{1}_{\rm cris}(A)$.
The statements about ${\rm H}^{i}_{\rm fppf}(A,\mu_{p^{n}})$ are well-known; we give a proof for the convenience of the reader. The Kummer sequence gives ${\rm H}^{0}_{\rm fppf}(A,\mu_{p^{n}})=0$, so in degree 0 there is nothing to prove. It also gives a natural isomorphism ${\rm H}^{1}_{\rm fppf}(A,\mu_{p^{n}})\cong{\rm{Pic}}(A)[p^{n}]$, which proves the statement in degree 1. The statement in degree 2 follows from the functoriality of the injective map ${\rm H}^{2}_{\rm fppf}(X,\mu_{p^{n}})\to{{\rm F}^{1}_{N}{}}{\rm H}_{\rm cris}^{2}(X)/p^{n}$ recalled in Remark 1.6. See also [YY, Cor. 5.7]. The last statement follows from the Kummer exact sequence. $\Box$
In general, the image of ${\bf H}_{\rm cris}^{2}(A/W)\to{\bf H}^{3}(A,\mu_{p^{n}})$ for an abelian variety $A$ can be non-trivial, for example when $A$ is the product of two supersingular elliptic curves [Gr25a, Prop. 5.2.6]. This shows that $[m]$ does not always act on ${\rm H}^{3}(A,\mu_{p^{n}})$ as multiplication by $m^{3}$.
2 Unipotent groups from Dieudonné modules
2.1 Unipotent groups attached to Dieudonné modules
Let $M$ and $N$ be Dieudonné modules. Let us write $M\otimes N:=M\otimes_{W}N$. The $W$-modules $M\otimes N$ and $\wedge^{2}M$ are F-crystals with the $\sigma$-linear action of $F$ given by $F(x\otimes y)=Fx\otimes Fy$ and $F(x\wedge y)=Fx\wedge Fy$. A natural isomorphism of F-crystals
$\wedge^{2}(M\oplus N)\cong(\wedge^{2}M)\oplus(\wedge^{2}N)\oplus(M\otimes N)$ (10)
gives a natural isomorphism of attached connected unipotent quasi-algebraic $k$-groups (see Definition 1.2):
${\bf U}_{\wedge^{2}(M\oplus N)}\cong{\bf U}_{\wedge^{2}M}\times{\bf U}_{\wedge^{2}N}\times{\bf U}_{M\otimes N}.$ (11)
Remark 2.1
If $N\subset M$ is an inclusion of Dieudonné modules, we write $M\wedge N$ for the image of $M\otimes N\subset M\otimes M$ in $\wedge^{2}M$.
Lemma 2.2
We have ${{\rm F}^{1}_{N}{}}(M)=VM$, ${{\rm F}^{1}_{N}{}}(M\otimes N)=VM\otimes N+M\otimes VN$, and ${{\rm F}^{1}_{N}{}}(\wedge^{2}M)=M\wedge VM$. In particular, the canonical map ${{\rm F}^{1}_{N}{}}(M\otimes M)\to{{\rm F}^{1}_{N}{}}(\wedge^{2}M)$ is surjective.
Proof. Since $F$ is injective, we have $Fx=py$ if and only if $x=Vy$, giving the first equality.
Let $m={\rm{rk}}_{W}(M)$ and $n={\rm{rk}}_{W}(N)$. We can reduce modulo $p$ and consider $k$-vector spaces $M_{1}=M/p$ and $N_{1}=N/p$ of dimensions $m$ and $n$, respectively. Let $m^{\prime}={\rm{dim}}_{k}(M_{1}[F])$ and $n^{\prime}={\rm{dim}}_{k}(N_{1}[F])$. Let $K$ be the kernel of $F$ acting on $M_{1}\otimes_{k}N_{1}$. It is clear that
$M_{1}[F]\otimes_{k}N_{1}+M_{1}\otimes_{k}N_{1}[F]\subset K.$
This subspace has dimension $m^{\prime}n+mn^{\prime}-m^{\prime}n^{\prime}=mn-(m-m^{\prime})(n-n^{\prime})$. The dimension of $F(M_{1}\otimes_{k}N_{1})=FM_{1}\otimes_{k}FN_{1}$ is $(m-m^{\prime})(n-n^{\prime})=mn-{\rm{dim}}(K)$, so the displayed inclusion is an equality. Since $M_{1}[F]=VM_{1}$ and $N_{1}[F]=VN_{1}$, we have $K=VM_{1}\otimes_{k}N_{1}+M_{1}\otimes_{k}VN_{1}$. The proof for $\wedge^{2}M$ is similar. $\Box$
Note that from ${{\rm F}^{1}_{N}{}}(M)=VM$ we deduce ${\bf U}_{M}=0$. Indeed, the map 1V1-V is surjective on $M$.
Remark 2.3
For any Dieudonné module $M$ we have
${\rm{dim}}({\bf U}_{\wedge^{2}M}[p])\geqslant a(a-1)/2,$
where $a={\rm{dim}}_{k}(M/(FM+VM))$ is the $a$-number of $M$. To prove this we note that F/p1F/p-1 sends ${{\rm F}^{1}_{N}{}}(\wedge^{2}M)=M\wedge VM$ into $M\wedge(FM+VM)$. This gives a surjection ${\bf U}_{\wedge^{2}M}\to{\bf G}_{a,k}^{a(a-1)/2}$. Comparing with (2) we see that this map is an isomorphism when $a=g$, which happens exactly when $M$ is the Dieudonné module of the superspecial abelian variety $E^{g}$, where $E$ is a supersingular elliptic curve.
We note that ${\rm{Hom}}_{W}(M,N)$ has a natural structure of a pro-algebraic group, which is isomorphic to ${\bf W}^{r_{1}r_{2}}$, where $r_{1}={\rm{rk}}_{W}(M)$ and $r_{2}={\rm{rk}}_{W}(N)$. This defines a quasi-algebraic group structure on ${\rm{Hom}}_{W}(M/p^{n},N/p^{n})$ and hence also on ${\rm{Hom}}_{\mathbb{E}}(M/p^{n},N/p^{n})$ which is the subgroup of ${\rm{Hom}}_{W}(M/p^{n},N/p^{n})$ invariant under the action of endomorphisms $F$ and $V$.
Proposition 2.4
Let $M$ and $N$ be Dieudonné modules. For every $n\geqslant 1$ there is a natural isomorphism of quasi-algebraic groups
${\bf U}_{M^{\vee}\otimes N}[p^{n}]\cong{\rm{Hom}}_{\mathbb{E}}(M/p^{n},N/p^{n})/\left({\rm{Hom}}_{\mathbb{E}}(M,N)/p^{n}\right).$
Proof. Applying Corollary 1.3 with $L=M^{\vee}\otimes N={\rm{Hom}}_{W}(M,N)$ we obtain that ${\bf U}_{L}[p^{n}]$ is naturally isomorphic to the cokernel of ${\bf L}^{F=p}/p^{n}\to{\bf H}_{L,n}$, where by definition we have ${\bf H}_{L,n}={\rm{Ker}}[F/p-1\colon{{\rm F}^{1}_{N}{}}({\bf L})/p^{n}\to{\bf L}/p^{n}]$.
The Frobenius $F$ acts on $L={\rm{Hom}}_{W}(M,N)$ by sending a homomorphism of $W$-modules $f\colon M\to N$ to $FfV$. We have $f\in L^{F=p}$ if and only if $pf=FfV$, which is equivalent to $Vf=fV$, and this implies $Ff=fF$. We conclude that $L^{F=p}={\rm{Hom}}_{\mathbb{E}}(M,N)$.
Next we note that ${{\rm F}^{1}_{N}{}}(L)$ is the set of $f\in L$ such that $p$ divides $FfV$. Thus $H_{L,n}=K_{L,n}/p^{n}{{\rm F}^{1}_{N}{}}(L)$, where
$K_{L,n}:=\{f\in L\ \text{such that}\ p|FfV(x)\ \text{and}\ p^{n}|FfV(x)/p-f(x)\ \text{for any}\ x\in M\}.$
The conjunction of these two conditions is equivalent to $p^{n+1}|FfV(x)-pf(x)$, thus
$K_{L,n}=\{f\in L\ \text{such that}\ p^{n+1}\ \text{divides}\ FfV(x)-pf(x)\ \text{for any}\ x\in M\}.$
Replacing $x$ by $Fx$ we deduce that $Ff\equiv fF\bmod p^{n}$. Similarly, applying $V$ to both sides we get $fV\equiv Vf\bmod p^{n}$. Thus reducing $f$ modulo $p^{n}$ gives a homomorphism
$\theta\colon K_{L,n}\to{\rm{Hom}}_{\mathbb{E}}(M/p^{n},N/p^{n}).$
By Corollary 1.3, it remains to prove that $\theta$ is surjective with ${\rm{Ker}}(\theta)=p^{n}{{\rm F}^{1}_{N}{}}(L)$.
It is clear that $p^{n}{{\rm F}^{1}_{N}{}}(L)\subset{\rm{Ker}}(\theta)$. Conversely, let $f\in{\rm{Ker}}(\theta)$. Then $f=p^{n}g$ for some $g\in{\rm{Hom}}_{W}(M,N)$ and we have that $p^{n+1}$ divides $p^{n}(FgV-pg)$. Hence $p|FgV$ so that $g\in{{\rm F}^{1}_{N}{}}(L)$. Thus ${\rm{Ker}}(\theta)=p^{n}{{\rm F}^{1}_{N}{}}(L)$.
To prove that $\theta$ is surjective, take any $\bar{f}\in{\rm{Hom}}_{\mathbb{E}}(M/p^{n},N/p^{n})$ and lift it to some $f\in{\rm{Hom}}_{W}(M,N)$. We need to find a $\phi\in{\rm{Hom}}_{W}(M,N)$ such that $f-p^{n}\phi\in K_{L,n}$, that is, such that $p^{n+1}|FfV-pf-p^{n}F\phi V$.
Since $Ff-fF\equiv 0\bmod p^{n}$, there is a $\sigma$-linear map $g\colon M\to N$ such that $Ff-fF=p^{n}g$. This implies $FfV-pf=p^{n}gV$. Likewise, there is a $\sigma^{-1}$-linear map $h\colon M\to N$ such that $fV-Vf=p^{n}h$. This implies $FfV-pf=p^{n}Fh$. The condition on $\phi$ can now be restated as the condition that $p$ divides $gV-F\phi V=Fh-F\phi V$.
Since $pM\subset VM\subset M$, we can find $W$-bases $e_{1},\ldots,e_{n}$ and $e^{\prime}_{1},\ldots,e^{\prime}_{n}$ of $M$ such that $Ve_{i}=e^{\prime}_{i}$ for $i=1,\ldots,m$ and $Ve_{i}=pe^{\prime}_{i}$ for $i=m+1,\ldots,n$, for some $m\leqslant n$. We now define $\phi$ so that $\phi(e^{\prime}_{i})=h(e_{i})$ for $i=1,\ldots,m$ and $\phi(e^{\prime}_{i})=0$ for $i=m+1,\ldots,n$. Then $(Fh-F\phi V)(e_{i})=0$ for $i=1,\ldots,m$. For $i=m+1,\ldots,n$ we know that $p$ divides $Ve_{i}$, hence $p$ divides $(gV-F\phi V)(e_{i})$. Thus the chosen $\phi$ satisfies the required condition. $\Box$
Corollary 2.5
Assume $p\neq 2$. Let $N$ be a Dieudonné module. For every $n\geqslant 1$ there is a natural isomorphism of quasi-algebraic groups
${\bf U}_{\wedge^{2}N}[p^{n}]\cong{\rm{Hom}}_{\mathbb{E}}(N^{\vee}/p^{n},N/p^{n})^{\rm skew}/\left({\rm{Hom}}_{\mathbb{E}}(N^{\vee},N)^{\rm skew}/p^{n}\right),$
where $\rm skew$} denotes the subgroup of anti-self-dual elements.
Proof. Let $\iota$ be the involution permuting the factors of $N\otimes N$. Since $p\neq 2$, we have isomorphisms
$\wedge^{2}(N/p^{n})\cong((N/p^{n})\otimes(N/p^{n}))^{\iota=-1}\cong{\rm{Hom}}_{W}(N^{\vee}/p^{n},N/p^{n})^{\rm skew}$
and compatible isomorphisms
$\wedge^{2}N\cong(N\otimes N)^{\iota=-1}\cong{\rm{Hom}}_{W}(N^{\vee},N)^{\rm skew}$
via the reduction mod $p$ maps. Passing to the attached unipotent quasi-algebraic groups we get an isomorphism ${\bf U}_{\wedge^{2}N}\cong{\bf U}_{(N\otimes N)^{\iota=1}}\cong({\bf U}_{N\otimes N})^{\iota=1}$. We thus obtain the desired isomorphism of unipotent quasi-algebraic groups by taking the $\iota$-anti-invariant subgroups in the isomorphism of Proposition 2.4 with $M=N$. $\Box$
2.2 Behaviour of ${\bf U}_{\wedge^{2}M}$ and ${\bf U}_{M\otimes N}$ under isogeny
Let $K=W[1/p]$ be the field of fractions of $W$. We call Dieudonné modules $M$ and $M^{\prime}$ isogenous if the $\mathbb{E}$}-modules $M\otimes_{W}K$ and $M^{\prime}\otimes_{W}K$ are isomorphic.
Proposition 2.6
The integers ${\rm{dim}}({\bf U}_{\wedge^{2}M})$ and ${\rm{dim}}({\bf U}_{M\otimes N})$ depend only on the isogeny classes of $M$ and $N$.
Proof. By (11) it is enough to prove the statement for ${\bf U}_{\wedge^{2}M}$.
Suppose that there is an exact sequence of $\mathbb{E}$}-modules
$0\to M\to M^{\prime}\to C\to 0,$ (12)
where $p^{n}C=0$ for some $n$. In particular, we have ${\rm{rk}}_{W}M={\rm{rk}}_{W}M^{\prime}$. We can consider (12) as an exact sequence in the abelian category $\mathcal{P}$} of pro-algebraic $k$-groups, with the maps respecting the action of $F$ and $V$ by $k$-endomorphisms. Applying the snake lemma to the commutative diagram obtained from the action of $V$ on $M$, $M^{\prime}$, $K$, we get an exact sequence of quasi-algebraic groups
$0\to C[V]\to M/VM\to M^{\prime}/VM^{\prime}\to C/VC\to 0.$
Counting dimensions we see that ${\rm{dim}}_{k}(M/VM)={\rm{dim}}_{k}(M^{\prime}/VM^{\prime})$. This implies
${\rm{dim}}_{k}(\wedge^{2}M/(M\wedge VM))={\rm{dim}}_{k}(\wedge^{2}M^{\prime}/(M^{\prime}\wedge VM^{\prime})).$
Defining the quasi-algebraic groups
$C_{1}=(M^{\prime}\wedge VM^{\prime})/(M\wedge VM),\quad C_{2}=\wedge^{2}M^{\prime}/\wedge^{2}M,$
we deduce that ${\rm{dim}}(C_{1})={\rm{dim}}(C_{2})$. We have the following commutative diagram of pro-algebraic groups
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\wedge VM\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F/p-1}$${M^{\prime}\wedge VM^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F/p-1}$${C_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\wedge^{2}M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\wedge^{2}M^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
where the rows are exact and the map $\phi$ is induced by two other vertical maps. The snake lemma gives an exact sequence
$0\to(\wedge^{2}M)^{F=p}\to(\wedge^{2}M^{\prime})^{F=p}\to{\rm{Ker}}(\phi)\to{\bf U}_{\wedge^{2}M}\to{\bf U}_{\wedge^{2}M^{\prime}}\to{\rm{Coker}}(\phi)\to 0.$
Now ${\rm{dim}}(C_{1})={\rm{dim}}(C_{2})$ implies that ${\rm{dim}}({\rm{Ker}}(\phi))={\rm{dim}}({\rm{Coker}}(\phi))$. The quotient of $(\wedge^{2}M^{\prime})^{F=p}$ by $(\wedge^{2}M)^{F=p}$ is a finite group. Thus counting dimensions of the terms of the exact sequence we obtain ${\rm{dim}}({\bf U}_{\wedge^{2}M})={\rm{dim}}({\bf U}_{\wedge^{2}M^{\prime}})$. $\Box$
Unlike the dimension, the $p$-exponent of ${\bf U}_{\wedge^{2}M}$ changes under isogeny. In certain cases, one can control this change.
Proposition 2.7
Let $M$ and $M^{\prime}$ be Dieudonné modules that fit into an exact sequence of $\mathbb{E}$}-modules
$0\to M\to M^{\prime}\to C\to 0.$
(i) If $VC=0$, then the kernel of the natural map ${\bf U}_{\wedge^{2}M}(k)\to{\bf U}_{\wedge^{2}M^{\prime}}(k)$ is annihilated by $p$.
(ii) If $FC=0$, then the cokernel of ${\bf U}_{\wedge^{2}M}(k)\to{\bf U}_{\wedge^{2}M^{\prime}}(k)$ is annihilated by $p$.
(iii) If $VC=FC=0$, then we have an exact sequence
$1\to{\bf G}_{a,k}^{n}\to{\bf U}_{\wedge^{2}M}\to{\bf U}_{\wedge^{2}M^{\prime}}\to{\bf G}_{a,k}^{n}\to 1$
for some $n\geqslant d(d-1)/2$, where $d={\rm{dim}}(C)$. In particular, the $p$-exponents of ${\bf U}_{\wedge^{2}M}$ and ${\bf U}_{\wedge^{2}M^{\prime}}$ differ at most by $1$.
Proof. We can think of $\wedge^{2}M$, $M\wedge VM$, $M^{\prime}\wedge VM^{\prime}$ as subcrystals of $\wedge^{2}M^{\prime}$.
In the notation of the previous proof we show that $p{\rm{Ker}}(\phi)=0$ if $VC=0$, and $p{\rm{Coker}}(\phi)=0$ if $FC=0$. This will prove (i) and (ii).
To prove the first statement, let $a\in M^{\prime}\wedge VM^{\prime}$ be such that $(F/p)(a)-a\in\wedge^{2}M$. We need to prove that $pa\in M\wedge VM$. The condition $VC=0$ implies that $VM^{\prime}\subset M$, and this implies $pM^{\prime}\wedge VM^{\prime}=VFM^{\prime}\wedge VM^{\prime}\subset\wedge^{2}M$. In particular, we have $pa\in\wedge^{2}M$, and hence $F(pa)\in p\wedge^{2}M$. This says that $pa\in{{\rm F}^{1}_{N}{}}(\wedge^{2}M)$. By Lemma 2.2 we have ${{\rm F}^{1}_{N}{}}(\wedge^{2}M)=M\wedge VM$, so $pa\in M\wedge VM$, as desired.
To prove the second statement, let $b\in\wedge^{2}M^{\prime}$. It is clear that $pb\in{{\rm F}^{1}_{N}{}}(\wedge^{2}M^{\prime})=M^{\prime}\wedge VM^{\prime}$. We can write pb=Fb(F/p1)(pb)pb=Fb-(F/p-1)(pb). The condition $FC=0$ implies that $FM^{\prime}\subset M$, and this implies $F(\wedge^{2}M^{\prime})\subset\wedge^{2}FM^{\prime}\subset\wedge^{2}M$. Hence the image of $pb$ in ${\rm{Coker}}(\phi)$ is zero, as required.
(iii) If $VC=FC=0$, then F/p1F/p-1 sends $M^{\prime}\wedge VM^{\prime}$ into $M\wedge M^{\prime}\subset\wedge^{2}M^{\prime}$. This gives $n\geqslant{\rm{dim}}_{k}\left(\wedge^{2}M^{\prime}/(M\wedge M^{\prime})\right)=d(d-1)/2$. $\Box$
2.3 The dimensions of ${\bf U}_{\wedge^{2}M}$ and ${\bf U}_{M\otimes N}$
We now calculate ${\rm{dim}}({\bf U}_{\wedge^{2}M})$ in terms of the Newton polygon of $M$, using only the classical Dieudonné theory and the invariance of ${\rm{dim}}({\bf U}_{\wedge^{2}M})$ under the isogeny of $M$.
Theorem 2.8
Let $M$ be a Dieudonné module with slopes $\lambda_{i}=c_{i}/(c_{i}+d_{i})$ of multiplicity $c_{i}+d_{i}$, where $c_{i}$ and $d_{i}$ are coprime non-negative integers, for $i=1,\ldots,n$ (allowing repetitions among the pairs $(c_{i},d_{i})$). Let $m_{i}\coloneqq\min(c_{i},d_{i})$. Then we have
${\rm{dim}}({\bf U}_{\wedge^{2}M})=\sum_{i=1}^{n}m_{i}(m_{i}-1)/2+\sum_{1\leq i<j\leq n}\min(c_{i}c_{j},d_{i}d_{j}).$
Proof. By Proposition 2.6, ${\rm{dim}}({\bf U}_{\wedge^{2}M})$ is an isogeny invariant of $M$, so we can replace $M$ by a direct sum of arbitrary isoclinic Dieudonné modules $M_{i}$ of rank $c_{i}+d_{i}$ and slope $c_{i}/(c_{i}+d_{i})$. Using (10) we obtain
${\rm{dim}}({\bf U}_{\wedge^{2}M})=\sum_{i=1}^{n}{\rm{dim}}({\bf U}_{\wedge^{2}M_{i}})+\sum_{1\leq i<j\leq n}{\rm{dim}}({\bf U}_{M_{i}\otimes M_{j}}).$ (13)
Theorem 2.8 is a consequence of Proposition 2.9 and Corollary 2.11 below.
Proposition 2.9
Let $M_{i}$ for $i=1,2$ be isoclinic Dieudonné modules of rank $c_{i}+d_{i}$ and slope $c_{i}/(c_{i}+d_{i})$, where $c_{i},d_{i}$ are coprime and non-negative. Then we have
${\rm{dim}}({\bf U}_{M_{1}\otimes M_{2}})=\min(c_{1}c_{2},d_{1}d_{2}).$
Proof. For integers $m\leqslant n$ we denote by $[m,n]$ the set of integers $m,m+1,\ldots,n$.
By Proposition 2.6 we can take $M_{1}=M_{c_{1},d_{1}}$. This is a free $W$-module with basis $e_{i}$ for $i\in[1,c_{1}+d_{1}]$ such that $Fe_{i}=e_{i+c_{1}}$ and $Ve_{i}=e_{i+d_{1}}$ with the convention that $e_{i+c_{1}+d_{1}}=pe_{i}$. Let us take $M_{2}=M_{c_{2},d_{2}}$ and denote the basis vectors of $M_{c_{2},d_{2}}$ by $f_{j}$ for $j\in[1,c_{2}+d_{2}]$. Write $H:=M_{c_{1},d_{1}}\otimes_{W}M_{c_{2},d_{2}}$. This is a free $W$-module with basis $v_{i,j}=e_{i}\otimes f_{j}$. Here
$(i,j)\in S_{0}\cup S_{1}\cup S_{2}\cup S_{3},$
where $S_{0}=[1,d_{1}]\times[1,d_{2}]$, $S_{1}=[d_{1}+1,c_{1}+d_{1}]\times[1,d_{2}]$, $S_{2}=[1,d_{1}]\times[d_{2}+1,c_{2}+d_{2}]$, $S_{3}=[d_{1}+1,c_{1}+d_{1}]\times[d_{2}+1,c_{2}+d_{2}]$. We note that $Fv_{i,j}=v_{i+c_{1},j+c_{2}}$ with the convention that $v_{i+c_{1}+d_{1},j}=v_{i,j+c_{2}+d_{2}}=pv_{i,j}$.
To prove the proposition, it is enough to show that ${\bf U}_{H}$ is the product of $\min(c_{1}c_{2},d_{1}d_{2})$ copies of ${\bf G}_{a,k}$. The proof of this consists of a series of claims.
Claim 1. ${{\rm F}^{1}_{N}{}}(H)$ is the $W$-submodule of $H$ with basis $pv_{i,j}$ for $(i,j)\in S_{0}$ and $v_{i,j}$ for $(i,j)\in S_{1}\cup S_{2}\cup S_{3}$. This is clear.
Claim 2. Let $\tau$ be an orbit of $F$ acting on $S$ by translation by $(c_{1},c_{2})$ subject to the convention mentioned above, and let $M_{\tau}$ be the $W$-submodule of $H$ generated by $v_{i,j}$ for $(i,j)\in\tau$. The cokernel of
$F/p-1\colon{{\rm F}^{1}_{N}{}}(M_{c_{1},d_{1}}\otimes M_{c_{2},d_{2}})\to M_{c_{1},d_{1}}\otimes M_{c_{2},d_{2}}$ (14)
is the direct sum of
${\bf U}_{\tau}:={\rm{Coker}}[F/p-1\colon{{\rm F}^{1}_{N}{}}(M_{\tau})\to M_{\tau}],$
where ${{\rm F}^{1}_{N}{}}(M_{\tau})=\{x\in M_{\tau}|p\ \text{divides}\ Fx\}$, over all orbits $\tau$. This is also clear.
We now fix an orbit $\tau$ of $F$ in $S$. We can think of $\tau$ as a cyclic graph with an orientation and refer to the elements of $\tau$ as vertices.
Claim 3. Assume that $d_{1}d_{2}\leqslant c_{1}c_{2}$. If $x,y\in\tau\cap S_{0}$, then there exists a vertex $z\in\tau\cap S_{3}$ between $x$ and $y$ in the direction from $x$ to $y$. Indeed, in the opposite case we have $z\in S_{1}\cup S_{2}$ for all $z\in\tau$ between $x$ and $y$. Let us show that this leads to a contradiction. Write $x=(x_{1},x_{2})$ and $y=(y_{1},y_{2})$. Suppose that there are exactly $t_{1}$ vertices of $S_{1}$ and $t_{2}$ vertices of $S_{2}$ in $\tau$ between $x$ and $y$ in the direction from $x$ to $y$. Using the description of the action of $F$ we obtain
$y_{1}-x_{1}$ $=(t_{1}+t_{2}+1)c_{1}-t_{1}(c_{1}+d_{1}),$
$y_{2}-x_{2}$ $=(t_{1}+t_{2}+1)c_{2}-t_{2}(c_{2}+d_{2}).$
Since $|y_{1}-x_{1}|<d_{1}$ and $|y_{2}-x_{2}|<d_{2}$, we get
$c_{1}(t_{2}+1)<d_{1}(t_{1}+1),\quad c_{2}(t_{1}+1)<d_{2}(t_{2}+1).$
Multiplying these inequalities we obtain $c_{1}c_{2}<d_{1}d_{2}$, which is a contradiction.
A similar proof shows that if $d_{1}d_{2}\geqslant c_{1}c_{2}$, then $\tau$ contains a vertex in $S_{0}$ between any two vertices in $S_{3}$.
Claim 4. Let us use the term chain for a connected and simply connected subgraph of $\tau$. A chain will be called relevant if all its vertices are in $S_{1}\cup S_{2}$ except the last one, which is in $S_{0}$; moreover, the vertex immediately preceeding the first vertex of the chain must be in $S_{3}$. The elements of relevant chains in $\tau$ will be called relevant. Let $L_{\tau}$ be the $W$-submodule of $M_{\tau}$ generated by $pv_{x}$ if $x\in\tau$ is relevant and by $v_{x}$ if $x\in\tau$ is not relevant. We claim that $L_{\tau}$ is contained in $(F/p-1)({{\rm F}^{1}_{N}{}}(M_{\tau}))$.
To prove this we define a linear map $\psi\colon L_{\tau}\to L_{\tau}$ by the rule: if $x$ is relevant, then $\psi(pv_{x})=pv_{x}$; if $x\in S_{0}$ is not relevant, then $\psi(v_{x})=pv_{x}$; and if $x\notin S_{0}$ is not relevant, then $\psi(v_{x})=v_{x}$. Next, define a $\sigma$-linear map $\phi\colon L_{\tau}\to L_{\tau}$ by the rule: if $x$ is relevant, then $\phi(pv_{x})=F(v_{x})$; if $x\in S_{0}$ is not relevant, then $\phi(v_{x})=F(v_{x})$; and if $x\notin S_{0}$ is not relevant, then $\phi(v_{x})=F(v_{x})/p$. It is easy to check that $\phi(L_{\tau})=L_{\tau}$, and by semilinear algebra this implies that $(\phi-\psi)(L_{\tau})=L_{\tau}$. A direct verification shows that $(\phi-\psi)(L_{\tau})$ is contained in $(F/p-1)({{\rm F}^{1}_{N}{}}(M_{\tau}))$, so the claim is proved.
Claim 5: ${\bf U}_{\tau}\simeq{\bf G}_{a,k}^{n}$, where $n$ is the number of relevant chains in $\tau$.
Let us prove this. The $k$-vector space $M_{\tau}/L_{\tau}$ has a basis consisting of the images of $v_{x}$ for all relevant $x\in\tau$. We can think of $M_{\tau}/L_{\tau}$ as a quasi-algebraic group given by a product of copies of ${\bf G}_{a,k}$ numbered by the relevant vertices in $\tau$.
Suppose that $v_{1},\ldots,v_{r}$ is a relevant chain in $\tau$. For any $i=1,\ldots,r-1$, we have $v_{i}\in{{\rm F}^{1}_{N}{}}(M_{\tau})$, thus $\sum_{i=1}^{r-1}(F/p-1)(Wv_{i})$ is contained in $(F/p-1)({{\rm F}^{1}_{N}{}}(M_{\tau}))$. This gives an injective map of quasi-algebraic groups ${\bf G}_{a,k}^{r-1}\to{\bf G}_{a,k}^{r}$, so its cokernel is isomorphic to ${\bf G}_{a,k}$. In fact, $(F/p-1)({{\rm F}^{1}_{N}{}}(M_{\tau}))$ is the sum of $L_{\tau}$ and the subgroups $\sum_{i=1}^{r-1}(F/p-1)(Wv_{i})$ for all relevant chains in $\tau$. Therefore, each relevant chain in $\tau$ contributes a copy of ${\bf G}_{a,k}$ to ${\bf U}_{\tau}$. This proves the claim.
End of proof of Proposition 2.9. From Claim 2 we easily deduce that the number of relevant chains in all orbits is $\min(c_{1}c_{2},d_{1}d_{2})$. Indeed, if $d_{1}d_{2}\leqslant c_{1}c_{2}$, then associating to a relevant chain to its last vertex defines a bijection between the set of relevant chains and $S_{0}$. Similarly, if $d_{1}d_{2}\geqslant c_{1}c_{2}$, then associating to a relevant chain the vertex immediately preceeding its first vertex defines a bijection between the set of relevant chains and $S_{3}$. Applying Claim 5 we prove the proposition. $\Box$
Remark 2.10
It is likely that this proposition can also be deduced from [Vas08, Examples 2.3.1, 2.3.2].
Corollary 2.11
Let $M$ be an isoclinic Dieudonné module of rank $c+d$ and slope $c/(c+d)$, where $c,d$ are coprime and non-negative. Then we have
${\rm{dim}}({\bf U}_{\wedge^{2}M})=m(m-1)/2,$
where $m=\min(c,d)$.
Proof. The $W$-module $\wedge^{2}M$ is the quotient of $M\otimes M$ by the $W$-submodule $K$ generated by $x\otimes x$, for $x\in M$, that is, in the notation of the previous proof, by $v_{i,i}$ for all $i$ and by $v_{i,j}+v_{j,i}$ for all $i\neq j$. By Lemma 2.2 the natural map ${{\rm F}^{1}_{N}{}}(M\otimes M)\to{{\rm F}^{1}_{N}{}}(\wedge^{2}M)$ is surjective, so we have a commutative diagram
${{{\rm F}^{1}_{N}{}}(M\otimes M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F/p-1}$${{{\rm F}^{1}_{N}{}}(\wedge^{2}M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F/p-1}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\wedge^{2}M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
Thus ${\bf U}_{\wedge^{2}M}(k)$ is the quotient of ${\bf U}_{M\otimes M}(k)$ by the image of $K$.
We now assume $d\leqslant c$, the proof in the opposite case being entirely similar. By the proof of Proposition 2.9, for $M=M_{c,d}$, the group ${\bf U}_{M\otimes M}$ is the product of copies of ${\bf G}_{a,k}$ numbered by the elements of $S_{0}=[1,d]\times[1,d]$. (Indeed, each relevant chain contributes a copy of ${\bf G}_{a,k}$, but for every $x\in S_{0}$ there is exactly one relevant chain that ends in $x$.) The involution swapping the factors of $M\otimes M$ acts on the complement to the diagonal in $1,d]\times[1,d]$ without fixed points, thus the image of $K$ in ${\bf U}_{M\otimes M}(k)\cong k^{d^{2}}$ is a vector subspace of dimension d+d(d1)/2d+d(d-1)/2. Hence ${\bf U}_{\wedge^{2}M}\cong{\bf G}_{a,k}^{d(d-1)/2}$. By Proposition 2.6 this proves the corollary. $\Box$
This finishes the proof of Theorem 2.8.
Corollary 2.12
Let $M$ be a Dieudonné module. Then ${\bf U}_{\wedge^{2}M}=0$ if and only if the slopes of $M$ are 0 (with any non-negative multiplicity), $1$ (with any non-negative multiplicity), and at most one slope $1/(1+n)$ or $n/(1+n)$ with multiplicity $1+n$.
Proof. The formula of Theorem 2.8 gives ${\rm{dim}}({\bf U}_{\wedge^{2}M})=0$ if and only if $m_{i}\in\{0,1\}$ for all $i$, and $m_{i}=1$ for at most one subscript $i$. This implies the statement. $\Box$
Corollary 2.13
Let $M$ and $N$ be a Dieudonné modules. Then ${\bf U}_{M\otimes N}=0$ if and only if at least one of $M$ and $N$ has all of its slopes 0 (with any non-negative multiplicity) and $1$ (with any non-negative multiplicity).
Proof. This immediately follows from Proposition 2.9. $\Box$
3 Abelian varieties
3.1 The dimension of ${\bf U}_{A}$
Using the results of the previous section it is now easy to deduce a formula for ${\rm{dim}}({\bf U}_{A})$, where $A$ is an abelian variety over $k$.
Theorem 3.1
Let $A$ be an abelian variety of dimension $g$ over $k$. Then we have
${\rm{dim}}({\bf U}_{A})=\frac{g(g-1)}{2}-\sum_{0\leqslant\lambda<1}(1-\lambda)m_{\lambda},$ (15)
where $m_{\lambda}$ is the multiplicity of slope $\lambda$ in ${\rm H}^{2}_{\rm cris}(A/W)$.
Proof. The contravariant Dieudonné module of $A$ is $M:={\rm H}^{1}_{\rm cris}(A/W)$, and there is a natural isomorphism ${\rm H}^{2}_{\rm cris}(A/W)\cong\wedge^{2}M$. Thus we have ${\bf U}_{A}={\bf U}_{\wedge^{2}M}$. By Proposition 2.6 we can assume without loss of generality that $M$ is the direct sum of isoclinic Dieudonné modules $M_{c,d}$ of rank $c+d$ and slope $c/(c+d)$, where $c,d$ are non-negative coprime integers. Since $M$ comes from an abelian variety, this decomposition is invariant under the symmetry swapping $c$ and $d$. Thus we can write
$M=M_{0}\oplus M_{1}\oplus\ldots\oplus M_{n},$
where $M_{0}\cong M_{1,1}^{\oplus m}$ for some $m\geqslant 0$, and $M_{i}=M_{c_{i},d_{i}}\oplus M_{d_{i},c_{i}}$ for $i=1,\ldots,n$, where $c_{i},d_{i}$ are non-negative and coprime with $c_{i}<d_{i}$. We have
${\rm{dim}}({\bf U}_{\wedge^{2}M})=\sum_{i=0}^{n}{\rm{dim}}({\bf U}_{\wedge^{2}M_{i}})+\sum_{0\leqslant i<j\leqslant n}{\rm{dim}}({\bf U}_{M_{i}\otimes M_{j}}).$ (16)
We can compute this using Proposition 2.9 and Corollary 2.11. Let us show by induction on $n$ that the answer agrees with formula (15).
This is true for $n=0$. Indeed, in this case $g=m$ and $\wedge^{2}M=\wedge^{2}M_{0}$ is isoclinic of slope 1, so (15) gives g(g1)/2g(g-1)/2. Applying Proposition 2.9 and Corollary 2.11 we check that (16) gives the same answer.
For $i\geqslant 1$, an application of Proposition 2.9 and Corollary 2.11 gives ${\rm{dim}}({\bf U}_{\wedge^{2}M_{i}})=c_{i}(c_{i}-1)+c_{i}d_{i}$. This agrees with formula (15) that gives
$\frac{(c_{i}+d_{i})(c_{i}+d_{i}-1)}{2}-\left(1-\frac{2c_{i}}{c_{i}+d_{i}}\right)\frac{(c_{i}+d_{i})(c_{i}+d_{i}-1)}{2}=c_{i}(c_{i}+d_{i}-1).$
Let us assume now that the two formulae agree for $M$ and show that they also agree for $M\oplus M_{n+1}$. Let $g=m+\sum_{i=1}^{n}(c_{i}+d_{i})$ be one half of the rank of $M$. Write $M_{n+1}=M_{c,d}\oplus M_{d,c}$, where $c<d$. By the above, it remains to check that
${\rm{dim}}({\bf U}_{M\otimes M_{n+1}})=g(c+d)-\sum_{0\leqslant\lambda<1}(1-\lambda)m_{\lambda},$
where $\lambda$ is a slope of $M\otimes M_{n+1}$. It is enough to check this for $M=M_{1,1}$ and for $M=M_{i}$, where $i=1,\ldots,n$.
Let $N=M_{1,1}\otimes(M_{c,d}\oplus M_{d,c})$. We need to check that
${\rm{dim}}({\bf U}_{N})=(c+d)-\sum_{0\leqslant\lambda<1}(1-\lambda)m_{\lambda},$
where $\lambda$ is a slope of $N$. Calculating slopes and their multiplicities, we obtain that the right hand side equals $2c$, and the same is obtained by applying Proposition 2.9 and Corollary 2.11.
Now let $N=(M_{a,b}\oplus M_{b,a})\otimes(M_{c,d}\oplus M_{d,c})$, where $a,b$ are coprime non-negative integers such that $a<b$. We need to check that
${\rm{dim}}({\bf U}_{N})=(a+b)(c+d)-\sum_{0\leqslant\lambda<1}(1-\lambda)m_{\lambda},$
where $\lambda$ is a slope of $N$. Without loss of generality we can assume that $a/b\geqslant c/d$. Calculating slopes and their multiplicities, we obtain that the right hand side of the last displayed formula gives $2c(a+b)$, and the same is obtained by applying Proposition 2.9 and Corollary 2.11. This finishes the proof. $\Box$
Remark 3.2
(i) Formula (15) implies that ${\rm{dim}}({\bf U}_{A})\leqslant g(g-1)/2$ with equality if and only if $g=1$ or $A$ is supersingular. Indeed, the slopes of ${\rm H}^{2}_{\rm cris}(A/W)$ are invariant under the reflection $\lambda\mapsto 2-\lambda$, so the equality is attained if and only if ${\rm H}^{2}_{\rm cris}(A/W)$ is isoclinic of slope $1$. For $g\geqslant 2$ this implies that ${\rm H}^{1}_{\rm cris}(A/W)$ is isoclinic of slope $1/2$.
(ii) We have ${\bf U}_{A}=0$ if and only if $A$ is ordinary (i.e., the slopes of ${\rm H}^{1}_{\rm cris}(A/W)$ are 0 and $1$, each with multiplicity $g$) or almost ordinary (i.e., the slopes of ${\rm H}^{1}_{\rm cris}(A/W)$ are 0 and $1$, each with multiplicity g1g-1, and $1/2$ with multiplicity $2$). This follows from Corollary 2.12 or from the proof of Theorem 3.1.
(iii) We have ${\bf U}_{A}={\bf G}_{a,k}$ if and only if the slopes of ${\rm H}^{1}_{\rm cris}(A/W)$ are 0 and $1$, each with multiplicity g2g-2, and $1/2$ with multiplicity 4. This can be deduced from the dimension formula of Theorem 2.8.
For details of the following statement see Section 3.2.
Corollary 3.3
The formal Brauer group of $A$ is smooth.
Proof. It is enough to check that the dimension of the formal Brauer group $\widehat{{\rm{Br}}}(A)$ equals the dimension of the tangent space to $\widehat{{\rm{Br}}}(A)$ at 0, which is ${\rm H}^{2}(A,{\cal O}_{A})$, see [AM77, Cor. II.2.4]. As we explain in Section 3.2 below, this equality is equivalent to formula (15). $\Box$
3.2 ${\bf U}_{A}$ and the formal Brauer group of AA
We now give a different computation of ${\rm{dim}}({\bf U}_{A})$, which exploits the relation between ${\bf U}_{A}$ and the formal Brauer group of $A$. For a smooth proper variety $X$ over $k$, the formal Brauer group $\widehat{\mathrm{Br}}(X)$ is the functor parameterising deformations of the trivial class in ${\rm{Br}}(X)$. For an Artinian $k$-algebra $R$ one defines
$\widehat{\mathrm{Br}}(X):={\rm{Ker}}[{\rm{Br}}(X_{R})\to{\rm{Br}}(X_{R_{\textnormal{red}}})],$ (17)
where $X_{R}=X\times_{k}{\rm{Spec}}(R)$. A detailed exposition of the formal Brauer group and its properties can be found in [AM77].
Proposition 3.4
Let $X$ be a smooth proper variety over $k$. The following properties hold.
$(1)$ If ${\rm{Pic}}_{X/k}$ is smooth, then $\widehat{\mathrm{Br}}(X)$ is representable by a connected formal group of finite type with tangent space ${\rm H}^{2}(X,\mathcal{O}_{X})$.
$(2)$ If $\widehat{{\rm{Br}}}(X)$ is representable and ${\rm H}_{\rm cris}^{3}(X/W)$ is torsion-free, then $\widehat{\mathrm{Br}}(X)$ is smooth.
$(3)$ If $\widehat{{\rm{Br}}}(X)$ is representable, then its covariant Dieudonné module is isomorphic to ${\rm H}^{2}(X,W\mathcal{O}_{X})$.
Proof. Parts (1) and (3) are proved in [AM77], see Proposition II.1.8, Corollaries II.2.4 and II.4.3. Part (2) is a consequence of work of Ekedahl [Eke85]. For a different proof, see [Gr25b, Corollary 3.17]. $\Box$
From now on we assume that $X$ is an abelian variety $\pi\colon A\to{\rm{Spec}}(k)$, so that the conclusions of Proposition 3.4 (1), (2), (3) are true. In particular, $\widehat{\mathrm{Br}}(A)$ is a smooth connected formal group of dimension $h^{0,2}(A)=g(g-1)/2$. Recall [Ha, (28.3.2)] that a smooth connected formal group of finite type is an extension of a smooth connected formal group of finite height by a smooth unipotent formal group, in a unique way. Let us denote by $\widehat{U}_{A}$ the unipotent part of $\widehat{\mathrm{Br}}(A)$, and by $\widehat{P}_{A}$ the maximal finite height quotient of $\widehat{\mathrm{Br}}(A)$. By construction, $\widehat{U}_{A}$ is the reduced formal subgroup of $\widehat{\mathrm{Br}}(A)[p^{n}]$ for large enough $n$.
By a theorem of Bragg and Olsson [BO21, Cor. 1.6] we know that $R^{2}\pi_{*}\mu_{p^{n},A}$ is represented by a group $k$-scheme. Let us denote by ${\mathcal{U}}_{A,n}$ its reduced connected component. The exact sequence
$1\to\mu_{p^{n},A}\to\mu_{p^{n+1},A}\to\mu_{p,A}\to 1$
[…]
Lemma 3.5
The formal group $\widehat{U}_{A}$ is canonically isomorphic to the formal completion of ${\mathcal{U}}_{A}$ at the identity.
Proof. A more general statement is proved in [BO21, version 1, Prop. 10.7]. We spell out the argument in our case. Recall that the formal completion functor maps a presheaf $\mathcal{F}$ over ${\rm{Spec}}(k)_{\rm fppf}$ to the functor
$\widehat{\mathcal{F}}:R\mapsto{\rm{Ker}}[\mathcal{F}(R)\to\mathcal{F}(R_{\textnormal{red}})]$
from the category of Artinian $k$-algebras to the category of abelian groups. Since the projection $R\to R_{\textnormal{red}}$ has a canonical retraction, the formal completion functor is exact. Pushing forward the Kummer sequence along $\pi:A\to{\rm{Spec}}(k)$ we obtain the short exact sequence of fppf sheaves
$0\to R^{1}\pi_{*}\mathbb{G}_{m,A}/p^{n}\to R^{2}\pi_{*}\mu_{p^{n},A}\to R^{2}\pi_{*}\mathbb{G}_{m,A}[p^{n}]\to 0,$
where $R^{1}\pi_{*}\mathbb{G}_{m,A}/p^{n}$ is representable by an étale group scheme, by the theory of the Picard scheme [CTS21, p. 121]. Therefore, taking the formal completion we get an isomorphism
$(R^{2}\pi_{*}\mu_{p^{n},A})^{\wedge}\cong(R^{2}\pi_{*}\mathbb{G}_{m,A}[p^{n}])^{\wedge}.$ (18)
Notice that the right-hand group is by definition $\widehat{\mathrm{Br}}(A)[p^{n}]$. Take $n$ large enough so that $\widehat{U}_{A}$ is the reduced formal subgroup of $\widehat{\mathrm{Br}}(A)[p^{n}]$, and ${\mathcal{U}}_{A}={\mathcal{U}}_{A,n}$. Then (18) gives a desired isomorphism $({\mathcal{U}}_{A})^{\wedge}\cong\widehat{U}_{A}$. $\Box$
By the Kummer sequence, the quasi-algebraic group $({\mathcal{U}}_{A})^{\rm perf}$ associated to ${\mathcal{U}}_{A}$ is isogenous to ${\bf U}_{A}$. Thus from the above discussion, we deduce the equalities:
${\rm{dim}}({\bf U}_{A})={\rm{dim}}(\widehat{U}_{A})={\rm{dim}}(\widehat{\mathrm{Br}}(A))-{\rm{dim}}(\widehat{P}_{A})=\frac{g(g-1)}{2}-{\rm{dim}}(\widehat{P}_{A}).$
It remains to compute ${\rm{dim}}(\widehat{P}_{A})$. By Proposition 3.4 (3), the covariant Dieudonné module of $\widehat{P}_{A}$ is isogenous to ${\rm H}^{2}(A,W\mathcal{O}_{A})$. The latter is isogenous to ${\rm H}^{2}_{\rm cris}(A/W)_{[0,1)}$ by the theory of the slope spectral sequence. It is well-known (see, e.g. [AM77, Proof of Cor. III.3.4]) that
${\rm{dim}}(\widehat{P}_{A})=\sum_{0\leqslant\lambda<1}m_{\lambda}(1-\lambda),$
where $\lambda$ is a slope of ${\rm H}^{2}_{\rm cris}(A/W)$ of multiplicity $m_{\lambda}$, see Lemma 3.6 below. This gives another proof of Theorem 3.1.
Lemma 3.6
Let $G$ be a $p$-divisible group. Let $M(G)$ be the covariant Dieudonné module of $G$. Then we have
${\rm{dim}}(G)=\sum_{\lambda}m_{\lambda}(1-\lambda),$
where $\lambda$ is a slope of $M(G)$ of multiplicity $m_{\lambda}$.
Proof. The dimension of a $p$-divisible group is an isogeny invariant, so we may assume that $M(G)$ is the simple isoclinic Dieudonné module $M_{c,d}$ as in the proof of Proposition 2.9. Thus $\lambda=c/(c+d)$ and $m_{\lambda}=c+d$. We can easily compute ${\rm{dim}}(G)={\rm{dim}}_{k}(M(G)/VM(G))=d=m_{\lambda}(1-\lambda)$. $\Box$
3.3 The $p$-exponent of ${\bf U}_{A}$ is at most g1g-1
In this section we prove the following theorem. The non-supersingular case uses Corollary 2.5 and thus requires the assumption $p\neq 2$.
Theorem 3.7
Let $A$ be an abelian variety of dimension $g\geqslant 1$ over an algebraically closed field $k$ of positive characteristic $p\neq 2$. Then $p^{g-1}{\bf U}_{A}=0$.
In Section 3.4 we show that the bound of Theorem 3.7 is best possible.
Corollary 3.8
Let $k_{0}$ be a field finitely generated over ${\mathbb{F}}_{p}$, where $p\neq 2$. Let $k$ be an algebraic closure of $k_{0}$. Let $A$ be an abelian variety over $k_{0}$ of dimension $g\geqslant 1$. Then the transcendental Brauer group ${\rm{Im}}[{\rm{Br}}(A)\to{\rm{Br}}(A_{k})]$ is a direct sum of a finite group and an abelian group annihilated by $p^{g-1}$.
Proof. This follows from Theorem 3.7 and a theorem of D’Addezio [D’A24, Thm. 5.2], see also [S24, Thm. 3.2]. $\Box$
Let $M={\rm H}^{1}_{\rm cris}(A/W)$ be the contravariant Dieudonné module of $A$. Recall that ${\bf U}_{A}={\bf U}_{\wedge^{2}M}$.
The non-supersingular case
Let $n\geqslant m\geqslant 1$. Precomposing a map $M^{\vee}/p^{m}\to M/p^{m}$ with the surjective map $M^{\vee}/p^{n}\to M^{\vee}/p^{m}$ and composing with the injective map $M/p^{m}\to M/p^{n}$ given by multiplication by $p^{n-m}$, we obtain an embedding
${\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{m},M/p^{m})\hookrightarrow{\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{n},M/p^{n}).$
Taking the anti-self-dual parts we get an embedding
${\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{m},M/p^{m})^{\rm skew}\hookrightarrow{\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{m+1},M/p^{m+1})^{\rm skew}.$
By Corollary 2.5 the $p$-exponent of ${\bf U}_{A}$ is the least positive integer $m$ such that
${\rm{dim}}\,{\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{m},M/p^{m})^{\rm skew}={\rm{dim}}\,{\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{n},M/p^{n})^{\rm skew}.$
Thus $\operatorname{\mathnormal{p}-exp}(A)\leqslant r$, where $r$ is the least positive integer such that for all $s\geqslant r$ we have
${\rm{dim}}\,{\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{r},M/p^{r})^{{\rm skew}}={\rm{dim}}\,{\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{s},M/p^{s})^{\rm skew}.$
For any Dieudonné modules $N$ and $N^{\prime}$ there exists an integer $f_{N,N^{\prime}}\geqslant 0$ such that the image of the reduction modulo $p^{m}$ map
${\rm{Hom}}_{\mathbb{E}}(N/p^{n},N^{\prime}/p^{n})\to{\rm{Hom}}_{\mathbb{E}}(N/p^{m},N^{\prime}/p^{m})$
is finite if and only if $n\geqslant m+f_{N,N^{\prime}}$, see [GV13, §6.1]. From [GV13, §6.1 (i), (ii)], see also [LNV13, Lemma 7.8], we obtain $r=f_{M^{\vee},M}$.
The abelian varieties $A$ and $A^{\vee}$ have the same Newton polygon that we denote by $\nu$. By [LNV13, Corollary 9.2] we have $f_{M^{\vee},M}\leqslant 2\nu(g)$. If $A$ is not supersingular, we have $\nu(g)<g/2$, hence $f_{M^{\vee},M}\leqslant g-1$. This finishes the proof in the non-supersingular case.
The supersingular case
Let $A$ be a supersingular abelian variety over $k$ of dimension $g\geqslant 1$. Let $E$ be a supersingular elliptic curve over $k$ and let $M_{E}$ be its Dieudonné module. We have $M_{E}\cong{\mathbb{E}}/(F-V){\mathbb{E}}$. Since $A$ is isogenous to $E^{g}$, we have an isomorphism of $\mathbb{E}$}-modules $M_{\mathbb{Q}}\cong M_{E,{\mathbb{Q}}}^{g}$. Recall that an abelian variety is called superspecial if it is isomorphic to $E^{n}$ for some $n\geqslant 1$. We call a Dieudonné module superspecial if it is isomorphic to $M_{E}^{n}$ for some $n\geqslant 1$. Recall that the $a$-number of $A$ is defined as
$a(A):={\rm{dim}}_{k}({\rm{Hom}}_{k}(\alpha_{p},A))={\rm{dim}}_{k}(M/(FM+VM)),$
hence the inequality $a(M)\leqslant g$.
Let us recall the basic theory of supersingular Dieudonné modules as developed in [Li89, §1]. By [Li89, Lemma 1.3] there exists a smallest superspecial Dieudonné module in $M_{\mathbb{Q}}\cong M_{E,{\mathbb{Q}}}^{g}$ containing $M$. Let us call it the superspecial envelope of $M$ and denote it by $S^{\prime}(M)$, using the notation of [Li89, §1]. The $W$-module $S^{\prime}(M)/M$ has finite length.
The $W$-module $S^{\prime}(M)$ is the direct sum of $g$ copies of $M_{E}=We\oplus Wf$, where $Fe=Ve=f$ and $Ff=Vf=pe$. Consider the filtration by the images of powers of Frobenius $F^{i}S^{\prime}(M)\subset S^{\prime}(M)$ for $i\geqslant 0$, where $F^{0}S^{\prime}(M)=S^{\prime}(M)$. Note that the same filtration is obtained by using $V$ instead of $F$. Thus each of $F$ and $V$ induces a ($\sigma$-linear or $\sigma^{-1}$-linear) bijective homomorphism of $g$-dimensional $k$-vector spaces:
$F^{i}S^{\prime}(M)/F^{i+1}S^{\prime}(M)\tilde{\longrightarrow}F^{i+1}S^{\prime}(M)/F^{i+2}S^{\prime}(M).$
This implies the injectivity of each of $F$ and $V$ on the graded factors of the induced filtration $M^{i}:=M\cap F^{i}S^{\prime}(M)$:
$M^{i}/M^{i+1}\hookrightarrow M^{i+1}/M^{i+2}.$
Let $n$ be the smallest integer such that $F^{n}S^{\prime}(M)\subset M$, which exists because $F$ is topologically nilpotent. Equivalently, $n$ is the smallest integer such that $M^{n}/M^{n+1}\simeq k^{g}$. (For all $i\geqslant n$ we have $M^{i}/M^{i+1}\simeq k^{g}$.)
Write $s_{i}={\rm{dim}}_{k}(M^{i}/M^{i+1})$, so $s_{n}=g$. Following [Li89, §1] we claim that if $s_{i}<g$, then
$F(M^{i}/M^{i+1})\neq V(M^{i}/M^{i+1}).$
In this case $M^{i+1}/M^{i+2}$ contains two distinct subspaces of dimension $s_{i}$, hence we have $s_{i}<s_{i+1}$ for $i\leqslant n-1$.
Let us prove the claim. By [Li89, Lemma 1.6], there is an element $v\in M$ whose $e_{i}$-coefficients reduced modulo $p$ are linearly independent over ${\mathbb{F}}_{p^{2}}$. (Equivalently, the image of $v$ in $M/M^{1}$ is not contained in $L\otimes_{{\mathbb{F}}_{p^{2}}}k$ for a ${\mathbb{F}}_{p^{2}}$-vector space $L$ of dimension at most g1g-1. If this is not the case, then $M$ is contained in a proper superspecial submodule of $S^{\prime}(M)$, which contradicts the fact that $S^{\prime}(M)$ is the superspecial envelope of $M$, see [Li89] for details.) If $s_{i}<g$ and $F(M^{i}/M^{i+1})=V(M^{i}/M^{i+1})$, then this is a proper $k$-vector subspace of the $k$-vector space $M^{i+1}/M^{i+2}$ of dimension at most g1g-1, which is invariant under $x\to x^{p^{2}}$, and so it is $L\otimes_{{\mathbb{F}}_{p^{2}}}k$ for some ${\mathbb{F}}_{p^{2}}$-vector space $L$. Then $F^{i+1}(v)\in L\otimes_{{\mathbb{F}}_{p^{2}}}k$, but this contradicts the definition of $v$.
Write $N=(F,V)M$ and consider the induced filtration $N^{i}:=N\cap F^{i}S^{\prime}(M)$. We have $(F,V)M\subset FS^{\prime}(M)=VS^{\prime}(M)$, so $N=N^{1}$. Define $t_{i}={\rm{dim}}_{k}(N^{i+1}/N^{i+2})$ for $i=0,\ldots,n-1$. The crucial inequality is $t_{i}\geqslant s_{i}+1$ for $i=0,\ldots,n-1$. This follows from the claim above. Indeed, for these values of $i$ we have two injective maps $M^{i}/M^{i+1}\hookrightarrow M^{i+1}/M^{i+2}$ (given by $F$ and $V$) whose images are different and contained in $N^{i+1}/N^{i+2}\subset M^{i+1}/M^{i+2}$.
We finally note that $F^{n}S^{\prime}(M)\subset M$ implies $F^{n+1}S^{\prime}(M)\subset(F,V)M$, hence $N^{n+1}=F^{n+1}S^{\prime}(M)=M^{n+1}$.
Since the $a$-number of $A$ is ${\rm{dim}}_{k}(M/(FM+VM))$, we obtain
$a(A)={\rm length}(M/M^{n+1})-{\rm length}(N/N^{n+1})=\sum_{i=0}^{n}s_{i}-\sum_{i=0}^{n-1}t_{i}=s_{n}-\sum_{i=0}^{n-1}(t_{i}-s_{i})\leqslant g-n.$
Let $A^{\prime}_{i}$ be an abelian variety with Dieudonné module $M^{i}$, for $i=0,\ldots,n$. In particular, $A^{\prime}_{0}=A$ and $A^{\prime}_{n}\simeq E^{g}$ is superspecial. Thus we get a chain of isogenies
$A=A^{\prime}_{0}\to A^{\prime}_{1}\to\ldots\to A^{\prime}_{n}\simeq E^{g},$
such that the kernel $K_{i}$ of $A^{\prime}_{i}\to A^{\prime}_{i+1}$ is $(\alpha_{p})^{s_{i}}$. It is easy to see that $a(A^{\prime}_{i+1})-a(A^{\prime}_{i})=t_{i}-s_{i}\geqslant 1$. It follows that $n\leqslant g-1$, and n=g1n=g-1 if and only if $a(A)=1$ and $a(A^{\prime}_{i})=i+1$ for $i=0,\ldots,g-1$. In this last case we have $s_{i}=i+1$ for $i=0,\ldots,g-1$, so that $K_{i}=\alpha(A^{\prime}_{i})$ is the group subscheme of $A^{\prime}_{i}$ generated by the images of morphisms $\alpha_{p}\to A^{\prime}_{i}$.
Applying previous considerations to the dual abelian variety, we obtain a chain of isogenies
$E^{g}\simeq A_{n}\to A_{n-1}\to\ldots\to A_{0}=A,$
whose successive kernels are $\alpha$-groups. Again, we have $n\leqslant g-1$, and n=g1n=g-1 if and only if $a(A)=1$ and $a(A_{i})=i+1$ for $i=0,\ldots,g-1$. In this case, the kernel of $A_{i}\to A_{i-1}$ is $(\alpha_{p})^{i}$, so in particular the kernel of $E^{g}\to A_{g-2}$ is $(\alpha_{p})^{g-1}$.
Proof of Theorem 3.7. By Proposition 2.7 and the fact that $p{\bf U}_{E^{g}}=0$, to prove that the $p$-exponent of ${\bf U}_{A}$ is at most g1g-1, we can assume that n=g1n=g-1, so that the kernel of $E^{g}\to A_{g-2}$ is $(\alpha_{p})^{g-1}$. Theorem 3.7 will follow if we show that ${\bf U}_{A_{g-2}}\to{\bf U}_{E^{g}}$ is the zero map.
We start with the case $g=2$.
Lemma 3.9
Let $E^{2}\to A$ be an isogeny with kernel $(\alpha_{p})^{n}$, where $n>0$. Then the induced map ${\bf U}_{A}\to{\bf U}_{E^{2}}$ is zero.
Proof. There is a commutative diagram with exact rows
${\wedge^{2}(M_{E^{2}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\bf U}_{E^{2}}(k)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${\wedge^{2}(M_{A})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\bf U}_{A}(k)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
The Dieudonné module $M_{E^{2}}=(M_{E})^{\oplus 2}$ has $W$-basis $e_{1},e_{2},f_{1},f_{2}$ such that $Fe_{i}=Ve_{i}=f_{i}$ and $Ff_{i}=Vf_{i}=pe_{i}$ for $i=1,2$. Thus ${{\rm F}^{1}_{N}{}}(\wedge^{2}(M_{E^{2}}))$ is generated by $pe_{1}\wedge e_{2}$, $f_{1}\wedge f_{2}$, and $e_{i}\wedge f_{j}$ for $i,j\in\{1,2\}$. Calculating the map
$F/p-1\colon{{\rm F}^{1}_{N}{}}(\wedge^{2}(M_{E^{2}}))\to\wedge^{2}(M_{E^{2}})$
shows that it is surjective onto ${{\rm F}^{1}_{N}{}}(\wedge^{2}(M_{E^{2}}))$, hence $\wedge^{2}(M_{E^{2}})\to{\bf U}_{E^{2}}(k)$ can be identified with the projection onto the 1-dimensional $k$-vector space $ke_{1}\wedge e_{2}$. By commutativity of the diagram, it is enough to show that the projection of $\wedge^{2}(M_{A})$ onto $ke_{1}\wedge ke_{2}$ is zero.
We have $n=2$ or $n=1$. In the first case, $M_{A}$ is the $W$-submodule of $M_{E^{2}}$ generated by $pe_{1},pe_{2},f_{1},f_{2}$, hence the projection of $\wedge^{2}(M_{A})$ onto $ke_{1}\wedge ke_{2}$ is zero. In the second case, $M_{A}$ is the $W$-submodule of $M_{E^{2}}$ generated by $pe_{1},pe_{2},f_{1},f_{2}$, and an element $a_{1}e_{1}+a_{2}e_{2}$, where at least one of $a_{1},a_{2}$ is non-zero modulo $p$. Once again the projection of $\wedge^{2}(M_{A})$ onto $ke_{1}\wedge ke_{2}$ is zero. $\Box$
Proposition 3.10
The map ${\bf U}_{A_{g-2}}\to{\bf U}_{E^{g}}$ is zero.
Proof. Write $E^{g}=\prod_{i=1}^{g}E^{(i)}$, where the $E^{(i)}\simeq E$ are isomorphic supersingular elliptic curves. It is known that
${\bf U}_{E^{g}}=\bigoplus_{i<j}p_{ij}^{*}({\bf U}_{E^{(i)}\times E^{(j)}})\cong({\bf G}_{k,a})^{g(g-1)/2},$
where $p_{ij}\colon E^{g}\to E^{(i)}\times E^{(j)}$ is the projection, see [YY, Prop. 4.14]. Thus it is enough to prove that the composition ${\bf U}_{A_{g-2}}\to{\bf U}_{E^{g}}\to{\bf U}_{E^{(i)}\times E^{(j)}}$ is zero. (The second map is induced by the natural morphism $E^{(i)}\times E^{(j)}\to E^{g}$.)
Let $B$ be the image of $E^{(i)}\times E^{(j)}$ in $A_{g-2}$. We have a commutative diagram
${E^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A}$${E^{(i)}\times E^{(j)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
It is enough to prove that the induced map ${\bf U}_{B}\to{\bf U}_{E^{(i)}\times E^{(j)}}$ is zero. The kernel of $E^{(i)}\times E^{(j)}\to B$ is the intersection of the kernel of $E^{g}\to A_{g-2}$, which is isomorphic to $(\alpha_{p})^{g-1}$, with $E^{(i)}\times E^{(j)}$. If this kernel was zero, we would get an injective map $(\alpha_{p})^{g-1}\hookrightarrow E^{g-2}$, which is not possible. Thus this kernel is $\alpha_{p}$ or $(\alpha_{p})^{2}$. It remains to apply Lemma 3.9 which says that ${\bf U}_{B}\to{\bf U}_{E^{(i)}\times E^{(j)}}$ is the zero map. $\Box$
3.4 The $p$-exponent of ${\bf U}_{A}$ can be g1g-1
The aim of this section is to prove the following statement.
Proposition 3.11
Assume that $p\geqslant 3$. For every $g\geqslant 1$ there exists a supersingular abelian variety $A$ of dimension $g$ over $k$ such that ${\bf U}_{A}$ is isogenous to $\prod_{i=1}^{g-1}{\bf W}_{i}$. In particular, the $p$-exponent of ${\bf U}_{A}$ is g1g-1.
Proof. By Corollary 2.5 it is enough to construct a Dieudonné module $M$ isoclinic of slope $1/2$, with a principal quasi-polarisation, and such that the connected component of the quasi-algebraic group ${\rm{End}}_{\mathbb{E}}(M/p^{n})^{\rm skew}$ is isogenous to $\prod_{i=1}^{g-1}{\bf W}_{i}$, for all large enough $n$. Here the superscript $\rm skew$}’ denotes the subgroup of anti-self-dual endomorphisms with respect to the principal quasi-polarisation. Indeed, let $S^{\prime}(M)$ be the superspecial envelope of $M$. We have $S^{\prime}(M)\cong M_{E^{g}}$, where $E$ is a supersingular elliptic curve over $k$. Dualising the embedding $M\to S^{\prime}(M)$, we obtain an isogeny of $p$-divisible groups $E^{g}[p^{\infty}]\to D$. Define $A$ as the quotient of $E^{g}$ by the kernel of this isogeny. Then ${\rm{dim}}(A)=g$ and the Dieudonné module of $A$ is $M$.
Following [NV07, Example 3.3], we consider $M={\mathbb{E}}/(F^{g}-V^{g}){\mathbb{E}}$. Thus $M$ is a free $W$-module of rank $2g$ with basis
$e_{i}=F^{i-1},\ \ i=1,\ldots,g,\quad e_{j}=V^{2g+1-j},\ \ j=g+1,\ldots,2g.$
We have $Fe_{i}=e_{i+1}$ for $i=1,\ldots,g$, and $Fe_{i}=pe_{i+1}$ for $i=g+1,\ldots,2g$ (with the convention that subscripts are considered as integers modulo $2g$.) By [LNV13, Lemma 5.8], the Dieudonné module $M$ is isoclinic of slope $1/2$.
The Dieudonné module $M$ has a principal quasi-polarisation, that is, a perfect, alternating, bilinear form
$\psi\colon M\otimes_{W}M\to W,$
such that $\psi(F(x),y)=\psi(x,V(y))^{\sigma}$, defined as follows. Let $\theta\in W$ be an invertible element such that $\sigma^{g}(\theta)=-\theta$. Then
$\psi(e_{i},e_{j})=0,\ \ i-j\neq g,\quad\psi(e_{i},e_{i+g})=\sigma^{i-1}(\theta),\ \ i=1,\ldots,g,$
is a principal quasi-polarisation. Write $B$ for the matrix of $\psi$ in the basis $e_{1},\ldots,e_{2g}$, so that $B_{ij}=\psi(e_{i},e_{j})$ for $i,j=1,\ldots,2g$.
Since $M$ is generated by $e_{1}$ as a $\mathbb{E}$}-module, there is at most one endomorphism $f\in{\rm{End}}_{\mathbb{E}}(M/p^{n})$ such that
$f(e_{1})=\sum_{i=1}^{2g}a_{i}e_{i},$
for given $a_{i}\in W_{n}=W/p^{n}$ for $i=1,\ldots,2g$, and there is exactly one such $f$ if and only if $F^{g}f(e_{1})=V^{g}f(e_{1})$. We have
$F^{g}f(e_{1})=\sum_{i=1}^{g}a_{i}^{\sigma^{g}}p^{i-1}e_{i+g}+\sum_{i=g+1}^{2g}a_{i}^{\sigma^{g}}p^{2g+1-i}e_{i-g},$
$V^{g}f(e_{1})=\sum_{i=1}^{g}a_{i}^{\sigma^{-g}}p^{i-1}e_{i+g}+\sum_{i=g+1}^{2g}a_{i}^{\sigma^{-g}}p^{2g+1-i}e_{i-g}.$
Thus $F^{g}f(e_{1})=V^{g}f(e_{1})$ is equivalent to the condition that the image of $a_{i}$ in $W_{n+1-i}$ is invariant under $\sigma^{2g}$, for $i=1,\ldots,2g$. Thus the connected component of ${\rm{End}}_{\mathbb{E}}(M)$ is isomorphic to
$\prod_{i=1}^{g}{\bf W}_{i-1}\times\prod_{i=g+1}^{2g}{\bf W}_{2g+1-i},$ (19)
where the $i$-th factor is a group subscheme of ${\bf W}_{n}$ with coordinate $a_{i}$.
Let $A$ be the $(2g\times 2g)$-matrix of $f$ in the basis $\{e_{1},\ldots,e_{2g}\}$, that is, $A_{ij}$ is the $e_{i}$-th coordinate of $f(e_{j})$. To $f$ we associate the map $\tilde{f}\in{\rm{Hom}}(M/p^{n},(M/p^{n})^{\vee})$ that sends $x$ to the linear form on $M/p^{n}$ whose value on $y$ is $\psi(Ax,y)$. The dual map $\tilde{f}^{\vee}\in{\rm{Hom}}(M/p^{n},(M/p^{n})^{\vee})$ sends $x$ to the linear form whose value on $y$ is $\psi(Ay,x)=-\psi(x,Ay)$. Thus $\tilde{f}^{\vee}=-f^{\vee}$ if and only if $A^{t}=BAB^{-1}$, or equivalently, $A=B^{-1}A^{t}B$. Since $f$ is an endomorphism of $\mathbb{E}$}-modules, this condition is equivalent to $A(e_{1})=B^{-1}A^{t}B(e_{1})$. We have $B(e_{1})=-\theta e_{g+1}$. Next, we use that $F^{i}(e_{1})=e_{1+i}$ for $i=1,\ldots,g$ and $V^{j}(e_{1})=e_{2g+1-j}$ for $j=1,\ldots,g$ to calculate
$A^{t}(e_{g+1})=a_{g+1}e_{1}+a_{g}^{\sigma}e_{2}+\ldots+a_{1}^{\sigma^{g}}e_{g+1}+a_{2g}^{\sigma^{-(g-1)}}e_{g+2}+\ldots+a_{g+2}^{\sigma^{-1}}e_{2g}.$
From this we obtain
$B^{-1}A^{t}B(e_{1})=a_{1}^{\sigma^{g}}e_{1}+\sum_{i=2}^{g}\frac{\theta}{\theta^{\sigma^{i-1}}}a_{2g+2-i}^{\sigma^{i-g-1}}e_{i}-\sum_{i=g+1}^{2g}\frac{\theta}{\theta^{\sigma^{i-g-1}}}a_{2g+2-i}^{\sigma^{i-g-1}}e_{i}.$
Thus the involution on ${\rm{End}}_{\mathbb{E}}(M/p^{n})$ that sends $A$ to $B^{-1}A^{t}B$ (which is indeed an involution because $B^{t}=-B^{-1}$) sends the $i$-th factor of (19) isomorphically onto the (2gi)(2g-i)-th factor, and acts on the $g+1$-th factor ${\bf W}_{g}$ as 1-1. Since $p\neq 2$, we conclude that the connected component of ${\rm{End}}_{\mathbb{E}}(M/p^{n})$ is isomorphic to $\prod_{i=1}^{g-1}{\bf W}_{i}$. $\Box$
3.5 Abelian threefolds
In this section we classify the isogeny classes of ${\bf U}_{A}$ for abelian threefolds $A$ over an algebraically closed field $k$ of characteristic $p\geqslant 3$. Recall that if $A$ is ordinary, then the slopes of the Newton polygon of $A$ are $(0,0,0,1,1,1)$, and if $A$ is supersingular, then all slopes are equal to $1/2$. There are three more cases, which we call almost ordinary if the slopes are $0,0,1/2,1/2,1,1$, almost supersingular if the slopes are $0,1/2,1/2,1/2,1/2,1$, and $1/3$-type if the slopes are $1/3,1/3,1/3,2/3,2/3,2/3$.
Let $\mathrm{BT}_{1}$ denote the category of $1$-truncated Barsotti–Tate groups, that is, finite commutative group $k$-schemes annihilated by $p$ such that $\mathrm{Ker}(F)=\mathrm{Im}(V)$ and $\mathrm{Ker}(V)=\mathrm{Im}(F)$. A $1$-truncated Barsotti–Tate group is said to be indecomposable if it is not a direct product of two non-trivial subgroup schemes.
Let us recall Kraft’s classification of $\mathrm{BT}_{1}$-group $k$-schemes [K75], see [O20, §5.1]. To every circular word $w$ composed of letters $F$ and $V$ one associates a $\mathrm{BT}_{1}$-group $G_{w}$ given by its covariant Dieudonné module $M_{w}$ which is an $\mathbb{E}$}-module of finite length annihilated by $p$, in particular it is a finite dimensional $k$-vector space. For example, the circular word $w=[FVF^{2}V^{2}]$ gives rise to the Dieudonné module $M_{w}$ with a $k$-basis $\{e_{1},e_{2},e_{3},e_{4},e_{5},e_{6}\}$ where $F$ and $V$ act as follows:
$e_{1}\xrightarrow{F}e_{2}\xleftarrow{V}e_{3}\xrightarrow{F}e_{4}\xrightarrow{F}e_{5}\xleftarrow{V}e_{6}\xleftarrow{V}e_{1}.$ (20)
This means that
$Fe_{1}=e_{2},\ Ve_{3}=e_{2},\ Fe_{3}=e_{4},\ Fe_{4}=e_{5},\ Ve_{6}=e_{5},\ Ve_{1}=e_{6};$ (21)
$Fe_{2}=Ve_{2}=Ve_{4}=Fe_{5}=Ve_{5}=Fe_{6}=0.$
The concatenation of two or more words corresponds to the direct sum of Dieudonné modules. We say that a word $w$ is indecomposable if it is not of the form $w=w^{\prime d}$ for some word $w^{\prime}$ and $d>1$.
Theorem 3.12
(Kraft) For every indecomposable circular word $w$, the $\mathrm{BT}_{1}$-group scheme $G_{w}$ is indecomposable. Every indecomposable $\mathrm{BT}_{1}$-group scheme is of the form $G_{w}$ for some indecomposable circular word $w$.
Proof. See, for example, [O20, Theorem 5.3]. $\Box$
Proposition 3.13
Let $A$ be an abelian variety over $k$ of dimension $g\geqslant 1$ with $a$-number 11 and $p$-rank 0. Then the following properties hold.
(i) There is an isomorphism of group $k$-schemes $A[p]\cong G_{w}$, where $w=[V^{g}F^{g}]$.
(ii) Assume, moreover, that $p\geqslant 3$. Then we have ${\rm{dim}}({\bf U}_{A}[p])=g-1$.
Proof. (i) Let $M$ be the covariant Dieudonné module of $A[p^{\infty}]$. Since the $a$-number of $A$ is 1, we have ${\rm{dim}}_{k}(M/(FM+VM))=1$. This implies that the $\mathbb{E}$}-module $M$ is generated by one element. Since the $p$-rank of $A$ is 0, both $F$ and $V$ are nilpotent on $M/pM$. Thus we can apply [LNV13, Lemma 5.7], which says that there is an isomorphism of $\mathbb{E}$}-modules $M\cong{\mathbb{E}}/f{\mathbb{E}}$, where
$f=a_{0}F^{g}+a_{1}F^{g-1}+\ldots+a_{g}+b_{1}V+\ldots+b_{g-1}V^{g-1}+b_{g}V^{g}\in{\mathbb{E}}$
with $a_{0},b_{g}\in W^{\times}$ and all other coefficients $a_{i},b_{j}$ in $pW$. Reducing modulo $p$, we obtain an $\mathbb{E}$}-module isomorphic to the Dieudonné module $M_{w}$ for $w=[V^{g}F^{g}]$.
(ii) By part (i), the $\mathrm{BT}_{1}$-group $k$-schemes $A[p]$ of abelian varieties $A$ of dimension $g$ with $a$-number 11 and $p$-rank 0 are isomorphic. By Corollary 2.5 the dimension of ${\bf U}_{A}[p]$ does not depend on the choice of $A$, and so can be calculated for the abelian variety with Dieudonné module ${\mathbb{E}}/(F^{g}-V^{g}){\mathbb{E}}$. Now (ii) follows from Proposition 3.11. $\Box$
Proposition 3.14
The following table lists all possibilities for the $\mathrm{BT}_{1}$-group $k$-scheme $A[p]\cong G_{w}$, where $A$ is an abelian threefold over $k$:
Newton polygon $a$ $w$
supersingular 33 $FV]^{3}$
supersingular 22 $FV][F^{2}V^{2}]$, $FVF^{2}V^{2}]$, $VFV^{2}F^{2}]$
supersingular 11 $F^{3}V^{3}]$
$1/3$ type 22 $F^{2}V][V^{2}F]$
$1/3$ type 11 $F^{3}V^{3}]$
almost supersingular 22 $F][V][FV]^{2}$
almost supersingular 11 $F][V][F^{2}V^{2}]$
almost ordinary 11 $F]^{2}[V]^{2}[FV]$
ordinary 0 $F]^{3}[V]^{3}$
Proof. The superspecial case ($a=3$), the ordinary case, the almost ordinary case, and the almost supersingular cases are clear. The supergeneral case and the $1/3$ type case with $a=1$ follow from Proposition 3.13 (i).
For the $1/3$ type case with $a=2$ and the supersingular case with $a=2$, the $p$-rank is 0. There are only four possibilities for the circular words that give rise to such group schemes:
$FVF^{2}V^{2}],\ \ [VFV^{2}F^{2}],\ \ [FV][F^{2}V^{2}],\ \ [F^{2}V][V^{2}F].$ (22)
The group scheme $G_{w}$ associated to $w=[F^{2}V][V^{2}F]$ is minimal in the sense of [O05, Definition, p. 1024]. In fact, we have $G_{w}\cong H_{1,2}[p]\oplus H_{2,1}[p]$, where $H_{c,d}$ is the $p$-divisible group with contravariant Dieudonné module $M_{c,d}$ as in the proof of Proposition 2.9. By minimality, $A[p]\cong G_{w}$ implies that $A[p^{\infty}]\cong H_{1,2}\oplus H_{2,1}$, so that the Newton polygon of $A$ is of $1/3$ type.
If $w$ is one of the three words in (22) other than $F^{2}V][V^{2}F]$, then
${\rm{dim}}({\rm{Hom}}_{\mathbb{E}}(M_{w},M_{w}^{\vee})^{{\rm skew}})=3.$ (23)
When $A[p]\cong G_{w}$, from Corollary 2.5 we obtain that ${\rm{dim}}({\bf U}_{A})\geqslant{\rm{dim}}({\bf U}_{A}[p])=3$. For abelian varieties $A$ of dimension $g$ we have ${\rm{dim}}({\bf U}_{A})\leqslant g(g-1)/2$ by (2), with equality if and only if $g=1$ or $A$ is supersingular (Remark 3.2 (i)). Thus (23) implies that our abelian threefold $A$ is supersingular.
The third named author proved (23) in his thesis, see [YY25, Section 4.4]. For the convenience of the reader let us give this calculation here.
Let $w=[FVF^{2}V^{2}]$. Let $e_{1}^{*},\ldots,e_{6}^{*}$ be the dual basis of the dual Dieudonné module $M_{w}^{\vee}$. Dualising (20) we obtain
$e_{1}^{*}\xleftarrow{V}e_{2}^{*}\xrightarrow{F}e_{3}^{*}\xleftarrow{V}e_{4}^{*}\xleftarrow{V}e_{5}^{*}\xrightarrow{F}e_{6}^{*}\xrightarrow{F}e_{1}^{*}.$ (24)
A linear map of $k$-vector spaces $f\colon M_{w}\to M_{w}^{\vee}$ is a map of Dieudonné modules if and only if it commutes with $F$ and $V$. Write $f(e_{j})=\sum a_{ij}e_{i}^{*}$ with $a_{ij}\in k$ such that $a_{ij}=-a_{ji}$ for all $i$ and $j$, in particular, $a_{ii}=0$. We consider subscripts as integers modulo 6.
Commuting with $F$ is equivalent to the conjunction of
(1) $a_{i+1,j+1}=a_{ij}^{\sigma}$, when $Fe_{j}=e_{j+1}$ and $Ve_{i+1}=e_{i}$;
(2) $a_{ij}=0$, when $Fe_{j}=0$ and $Ve_{i+1}=e_{i}$.
Commuting with $V$ is equivalent to the conjunction of
(3) $a_{i+1,j+1}=a_{i,j}^{\sigma}$, when $Ve_{j+1}=e_{j}$ and $Fe_{i}=e_{i+1}$;
(4) $a_{ij}=0$, when $Ve_{j}=0$ and $Fe_{i-1}=e_{i}$.
In our case $(i,j)$ satisfies condition (1) for $i=2,5,6$ and $j=1,3,4$, condition (2) for $i,j\in\{2,5,6\}$, condition (3) for $i=1,3,4$ and $j=2,5,6$, and condition (4) for $i,j\in\{2,4,5\}$. The set of pairs $(i,j)$ with $i\neq j$ is the disjoint union of chains where we have an arrow $(i,j)\mapsto(i+1,j+1)$ if $a_{i+1,j+1}=a_{ij}^{\sigma}$. Counting the chains that do not contain a pair $(i,j)$ satisfying (2) or (4), we see that there are exactly six such chains. Taking skew-symmetry of $(a_{ij})$ into account, we deduce that ${\rm{dim}}_{k}({\rm{Hom}}(M_{w},M_{w}^{\vee})^{{\rm skew}})=3$.
The case $w=[VFV^{2}F^{2}]$ is entirely similar.
It remains to deal with the case $w=[FV][F^{2}V^{2}]$. Let $M$ and $N$ be the Dieudonné modules given by $[FV]$ and $F^{2}V^{2}]$, respectively. Thus $M_{w}=M\oplus N$, and we have
${\rm{Hom}}_{\mathbb{E}}(M_{w},M_{w}^{\vee})^{{\rm skew}}\cong{\rm{Hom}}_{\mathbb{E}}(M,M^{\vee})^{{\rm skew}}\oplus{\rm{Hom}}_{\mathbb{E}}(N,N^{\vee})^{{\rm skew}}\oplus{\rm{Hom}}_{\mathbb{E}}(M,N^{\vee}).$
Since ${\bf U}_{\wedge^{2}(M)}=0$, we have ${\rm{Hom}}_{\mathbb{E}}(M,M^{\vee})^{{\rm skew}}=0$. Next, ${\bf U}_{\wedge^{2}(N)}\cong{\bf G}_{a,k}$ implies that ${\rm{Hom}}_{\mathbb{E}}(N,N^{\vee})^{{\rm skew}}$ has dimension 1. Noting that $N$ is self-dual, it remains to check that the dimension of ${\rm{Hom}}_{\mathbb{E}}(M,N)$ is 2. Since $M={\mathbb{E}}/(F-V){\mathbb{E}}$, this dimension is equal to the dimension of $N^{F=V}$. The Dieudonné module $N$ is a $k$-vector space with basis $e_{1},e_{2},e_{3},e_{4}$ with action of $F$ and $V$ given by
$e_{1}\xrightarrow{F}e_{2}\xrightarrow{F}e_{3}\xleftarrow{V}e_{4}\xleftarrow{V}e_{1}.$
Write $u=a_{1}e_{1}+a_{2}e_{2}+a_{3}e_{3}+a_{4}e_{4}\in N$. Then $F(u)=a_{1}^{\sigma}e_{2}+a_{2}^{\sigma}e_{3}$ and $V(u)=a_{1}^{\sigma^{-1}}e_{4}+a_{4}^{\sigma^{-1}}e_{3}$. Hence $N^{F=V}$ is given by $a_{1}=0$ and $a_{4}=a_{2}^{\sigma^{2}}$. Thus $a_{2}$ and $a_{3}$ are free variables, so the dimension of $N^{F=V}$ is 2. This finishes the proof. $\Box$
Proposition 3.15
Let $A$ be an abelian threefold over an algebraically closed field $k$ of characteristic $p\geqslant 3$. The following table lists all possible isogeny classes of ${\bf U}_{A}$:
Newton polygon $a$ ${\rm{dim}}({\bf U}_{A}[p])$ ${\rm{dim}}({\bf U}_{A})$ ${\bf U}_{A}$ up to isogeny
supersingular 33 33 33 $({\bf G}_{a,k})^{3}$
supersingular 22 33 33 $({\bf G}_{a,k})^{3}$
supersingular 11 22 33 ${\bf G}_{a,k}\times{\bf W}_{2}$
$1/3$ type 22 22 22 $({\bf G}_{a,k})^{2}$
$1/3$ type 11 22 22 $({\bf G}_{a,k})^{2}$
almost supersingular 22 11 11 ${\bf G}_{a,k}$
almost supersingular 11 11 11 ${\bf G}_{a,k}$
almost ordinary 11 0 0 0
ordinary 0 0 0 0
In particular, the isogeny class of ${\bf U}_{A}$ is uniquely determined by the Newton polygon and the $a$-number of $A$. We have ${\bf U}_{A}\simeq({\bf G}_{a,k})^{n}$, where $0\leqslant n\leqslant 3$, unless $A$ is supergeneral, in which case ${\bf U}_{A}$ is isogenous to ${\bf G}_{a,k}\times{\bf W}_{2}$.
Proof. The superspecial case is clear. Using the dimension formula (2), the supergeneral case and the $1/3$ type case with $a=1$ follow from Proposition 3.13 (ii). We dealt with the supersingular case with $a=2$ in the proof of Proposition 3.14 which shows that $pU_{A}=0$ in this case.
For the $1/3$ type case with $a=2$, we have $A[p]\cong G_{w}$ where $w=[F^{2}V][V^{2}F]$. As was observed in the proof of Proposition 3.14, this implies that $A[p^{\infty}]\cong H_{1,2}\oplus H_{2,1}$. In this case we have $pU_{A}=0$, as follows from the proofs of Proposition 2.9 and Corollary 2.11.
The remaining cases were handled in Remark 3.2 (ii) and (iii). $\Box$
4 Ekedahl–Oort types and ${\rm{dim}}({\bf U}_{A}[p])$
Let $A$ be an abelian variety of dimension $g$ over an algebraically closed field $k$ of characteristic $p\geqslant 3$. In this section we compute ${\rm{dim}}({\bf U}_{A}[p])$ when $A$ is principally polarised in terms of the Ekedahl–Oort type of $A$.
In [O01], Ekedahl and Oort stratified the moduli space of principally polarised abelian varieties of dimension $g$ over $k$ according to the isomorphism class of $A[p]$. In total there are $2^{g}$ such isomorphism classes; they are numbered by the Ekedahl–Oort types. An Ekedahl–Oort type is a non-decreasing function
$\varphi\colon\{0,1,\ldots,g\}\to\{0,1,\ldots,g\},\quad\varphi(0)=0,$
such that $\varphi(i)-\varphi(i-1)$ is 0 or 1. Given a subset $P\subset\{1,\ldots,g\}$ we define $\varphi_{P}(i)=\varphi_{P}(i-1)+1$ if $i\in P$, and $\varphi_{P}(i)=\varphi_{P}(i-1)$ otherwise. It is clear that every Ekedahl–Oort type can be written as $\varphi_{P}$ for some subset $P\subset\{1,\ldots,g\}$. The correspondence between Ekedahl–Oort types and isomorphism classes of $A[p]$ is recalled in Theorem 4.3. Note that the $a$-number of $A$ is g|P|g-|P|.
Theorem 4.1
Let $A$ be a principally polarised abelian variety of dimension $g$ over an algebraically closed field of characteristic $p\geqslant 3$. If the Ekedahl–Oort type of $A$ is $\varphi_{P}$, where $P=\{m_{1}<\ldots<m_{g-a}\}$, where $a$ is the $a$-number of $A$, then
${\rm{dim}}({\bf U}_{A}[p])=\frac{a(a-1)}{2}+\sum_{i=1}^{g-a}(m_{i}-i).$
When $P=\emptyset$, the sum is understood to be 0.
Example 4.2
Let $A$ be a principally polarised abelian variety.
(1) Suppose that $A$ is supersingular. Then we have ${\rm{dim}}({\bf U}_{A})=g(g-1)/2$. There are two extreme cases:
  • AA is superspecial if and only if $a=g$, which is equivalent to the condition that $\varphi$ is the zero function (that is, $P=\emptyset$). Theorem 4.1 gives ${\rm{dim}}({\bf U}_{A}[p])=g(g-1)/2$, so that $p{\bf U}_{A}=0$. This is already proved in [$Y$, Proposition 4.4].
  • AA is supergeneral if and only if $a=1$. Then $P=\{2,\ldots,g\}$ and $\varphi(i)=i-1$ for $i\geqslant 1$. (This can be deduced from [O01, (9.7)] using that the $p$-rank is 0.) Theorem 4.1 says that ${\rm{dim}}({\bf U}_{A}[p])=g-1$.
(2) Suppose that $A$ is ordinary. This is equivalent to $a=0$, and thus to $\varphi(i)=i$ for all $i=0,1,\ldots,g$.
We now recall the construction of covariant self-dual Dieudonné modules associated to Ekedahl–Oort types [O01, §9]. Let $P=\{m_{1}<...<m_{h}\}$ be a subset of $\{1,2,...,g\}$ of size h=gah=g-a, and let $\varphi=\varphi_{P}$. Extend $\varphi$ to a non-decreasing function
$\psi\colon\{0,1,\ldots,2g\}\to\{0,1,\ldots,g\},\quad\psi(0)=0,\ \psi(2g)=g,$
by setting $\psi(i)=\varphi(2g-i)+i-g$ for i=g+1,,2gi=g+1,...,2g. Let $m_{1}<\ldots<m_{g}$ be the increasing sequence of all integers $1\leqslant i\leqslant 2g$ such that $\psi(i)=\psi(i-1)+1$. Likewise, let $n_{1}>\ldots>n_{g}$ be the decreasing sequence of all integers $1\leqslant j\leqslant 2g$ such that $\psi(j)=\psi(j-1)$. In particular, $\{1,...,g\}$ is the disjoint union of $P=\{m_{1},\ldots,m_{h}\}$ and its complement $Q=\{n_{g},\ldots,n_{h+1}\}$.
Let $M_{\varphi}$ be a vector space over ${\mathbb{F}}_{p}$ with basis $Z_{1},\ldots,Z_{2g}$. Let us write ${\cal Z}_{1}:={\rm Span}(Z_{1},\ldots,Z_{g})$ and ${\cal Z}_{2}:={\rm Span}(Z_{g+1},\ldots,Z_{2g})$. We rename the set of basis vectors
$Z_{m_{i}}=X_{i},\quad Z_{n_{i}}=Y_{i},\ \ \text{where}\ \ i=1,\ldots,g,$
and write ${\cal X}:={\rm Span}(X_{i})$, ${\cal Y}:={\rm Span}(Y_{j})$. The ${\mathbb{F}}_{p}$-vector space $M_{\varphi}$ has an alternating bilinear form $\langle x,y\rangle$ which is zero on $\cal X$} and $\cal Y$}, and satisfies $\langle X_{i},Y_{j}\rangle=\delta_{ij}$, where $\delta_{ij}$ is Kronecker’s delta. We have $m_{i}+n_{i}=2g+1$ for all $i=1,\ldots,g$. Using this one checks that ${\cal Z}_{1}$ and ${\cal Z}_{2}$ are maximal hyperbolic subspaces, alongside with $\cal X$} and $\cal Y$}.
Define $\mathcal{F}\in{\rm{End}}_{{\mathbb{F}}_{p}}(M_{\varphi})$ as follows:
$\mathcal{F}(X_{i})=Z_{i},\quad\mathcal{F}(Y_{i})=0.$
We have $\langle\mathcal{F}x,\mathcal{F}y\rangle=0$ for all $x,y\in M_{\varphi}$, because ${\rm{Im}}(\mathcal{F})={\cal Z}_{1}$ which is maximal hyperbolic. Define $\mathcal{V}\in{\rm{End}}_{{\mathbb{F}}_{p}}(M_{\varphi})$ so that
$\langle\mathcal{F}x,y\rangle=\langle x,\mathcal{V}y\rangle$ (25)
for all $x,y\in M_{\varphi}$. By the non-degeneracy of the form we have $\mathcal{V}\mathcal{F}=0$. Using this non-degeneracy and the fact that ${\rm{Ker}}(\mathcal{F})={\cal Y}$, we see that ${\rm{Im}}(\mathcal{V})={\cal Y}$ and ${\rm{Ker}}(\mathcal{V})={\cal Z}_{1}$. We deduce that $\langle\mathcal{V}x,\mathcal{V}y\rangle=0$ for all $x,y\in M_{\varphi}$. This implies $\mathcal{F}\mathcal{V}=0$.
Theorem 4.3
(Oort) Let $k$ be an algebraically closed field of positive characteristic $p\neq 2$. Let $A$ be a principally polarised abelian variety over $k$ of Ekedahl–Oort type $\varphi$. Then the covariant Dieudonné module $M$ of $A[p]$ is $M\cong M_{\varphi}\otimes k$ with the natural $\sigma$-linear action of $\mathcal{F}$ and the natural $\sigma^{-1}$-linear action of $\mathcal{V}$. The Dieudonné module $M$ is self-dual with respect to the alternating bilinear form $\langle x,y\rangle$.
Proof. This is [O01, Theorem 9.4]. $\square$
Let $E_{i,j}$ be the $(g\times g)$-matrix with $(i,j)$-entry 1 and all other entries 0. Define
$U=\sum_{i=1}^{h}E_{i,m_{i}},\quad V=\sum_{j=h+1}^{g}E_{j,n_{j}}.$
The matrix $U$ encodes that $\mathcal{F}(X_{m_{i}})=Z_{m_{i}}=X_{i}$, whereas $V$ encodes that $\mathcal{F}(X_{n_{i}})=Z_{n_{i}}=Y_{i}$. Thus $U+V$ is the permutation matrix of the permutation of $\{1,\ldots,g\}$ defined on $P$ as the unique order-preserving function $P\to\{1,\ldots,h\}$ (the inverse of $i\mapsto m_{i}$) and on $Q$ as the unique order-inversing function $Q\to\{h+1,\ldots,g\}$ (the inverse of $j\mapsto n_{j}$).
Lemma 4.4
The group of self-dual endomorphisms of the self-dual Dieudonné module $M$ is isomorphic to the additive group of pairs of $(g\times g)$-matrices $(A,C)$ with entries in $k$ such that
$C^{t}=-C,\quad VA=C^{(p)}U+A^{t(p)}V,\quad UA=A^{(p)}U,$ (26)
where superscript $(p)$ means that every entry of the matrix is raised to the power $p$.
Proof. We use the basis $\mathcal{B}=\{X_{1},\ldots,X_{g},Y_{1},\ldots,Y_{g}\}$. In this basis the matrix of the form $\langle x,y\rangle$ is
$\left(\begin{matrix}0&I\\ -I&0\end{matrix}\right).$
The linear action of $\mathcal{F}$ and $\mathcal{V}$ on $M_{\varphi}$ in the basis $\mathcal{B}$ is given by the matrices
$\mathcal{[}\mathcal{F}]=\left(\begin{matrix}U&0\\ V&0\end{matrix}\right),\quad\quad[\mathcal{V}]=\left(\begin{matrix}0&0\\ -V^{t}&U^{t}\end{matrix}\right).$
Here $\mathcal{V}]$ is obtained from $\mathcal{F}$] using (25).
Let $f$ be an endomorphism of the self-dual Dieudonné module $M$. We have $f=f^{\vee}$ if and only if the matrix of $f$ in the basis $\mathcal{B}$ is
$f]=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)$
where $D=A^{t}$, $B^{t}=-B$, $C=-C^{t}$. The condition $\mathcal{F}f=f^{\sigma}\mathcal{F}$ is equivalent to
$UA=A^{(p)}U+B^{(p)}V,\quad VA=C^{(p)}U+A^{t(p)}V,\quad UB=VB=0,$
Since $U+V$ is invertible, we have $B=0$. This implies the lemma. $\Box$
As an example, suppose that $P=\{1,\ldots,g\}$ so that $Q=\emptyset$. Then $U=I$ and $V=0$. It follows that $C=0$ and $A$ is an arbitrary element of ${\rm Mat}_{g}({\mathbb{F}}_{p})$. This is the case for ordinary abelian varieties. Now suppose that $P=\emptyset$. Then $U=0$ and $V$ is the matrix of the permutation $i\mapsto g+1-i$. Then $C$ is an arbitrary skew-symmetric matrix with entries in $k$. Since $p\neq 2$, this gives a vector space over $k$ of dimension g(g1)/2g(g-1)/2. The condition on $A$ gives a finite-dimensional vector space over ${\mathbb{F}}_{p}$. This is the case for superspecial abelian varieties. Both results agree with Theorem 4.1.
Proof of Theorem 4.1. By Corollary 2.5 it is enough to prove the following statement.
($\star$) The subgroup $S\subset{\rm Mat}_{g}(k)$, defined by the property that for every $A\in S$ there is a matrix $C\in{\rm Mat}_{g}(k)$ with $(A,C)$ a solution of (26), is isomorphic to a direct sum of a finite-dimensional ${\mathbb{F}}_{p}$-vector space and a $k$-vector space of dimension $\sum_{i=1}^{h}(m_{i}-i)$.
Indeed, by ($\star$), for every $A\in S$ there is a matrix $C$ such that $(A,C)$ satisfies (26). If $(A,C+C^{\prime})$ is a solution of (26), then $C^{\prime}U=0$ which says that the $h$ columns of the skew-symmetric matrix $C^{\prime}$ corresponding to the elements of $P$ are zero. Thus $C$ is uniquely determined up to addition of an element of a $k$-vector subspace of ${\rm Mat}_{g}(k)$ of dimension (gh)(gh1)/2(g-h)(g-h-1)/2, so ($\star$) implies the theorem.
Lemma 4.5
The subgroup $S\subset{\rm Mat}_{g}(k)$ is given by the following conditions:
(1) $a_{m_{i},m_{j}}=a_{i,j}^{p}$ for $i,j=1,\ldots,h$;
(2) $a_{n_{j},n_{i}}=a_{i,j}^{p}$ for $i,j=h+1,\ldots,g$;
(3) if $i\geqslant h+1$, $j\leqslant h$, then $a_{i,j}=0$;
(4) if $i\in P$, $j\in Q$, then $a_{i,j}=0$.
Proof. The $i$-th row of $UA$ is the $m_{i}$-th row of $A$, for $i=1,\ldots,h$, and all other rows are zero. The $m_{j}$-th column of $A^{(p)}U$ is the $j$-th column of $A^{(p)}$, for $j=1,\ldots,h$, and all other columns are zero. This implies that $UA=A^{(p)}U$ is equivalent to the conjunction of conditions (1), (3), (4).
The $m_{j}$-th column of $C^{(p)}U$ is the $j$-th column of $C^{(p)}$, for $j=1,\ldots,h$, and all other columns are zero. The $n_{j}$-th column of $A^{t(p)}V$ is the $j$-th row of $A^{(p)}$, for $j=h+1,\ldots,g$, and all other columns are zero. The $i$-th row of $VA$ is the $n_{i}$-th row of $A$, for $i=h+1,\ldots,g$, and all other rows are zero. From this one sees that the conditions imposed on $A$ by the equation $VA=C^{(p)}U+A^{t(p)}V$ are equivalent to the conjunction of (3) and (2). $\square$
To prove ($\star$) we consider a directed graph $\mathcal{G}_{\varphi}$, whose vertices are pairs $(i,j)$, for $i,j=1,\ldots,g$, and the directed edges encode the relations of Lemma 4.5. Namely, the edges of $\mathcal{G}_{\varphi}$ are of the following types (loops are allowed):
(m) for $i,j=1,\ldots,h$ there is an edge from $(i,j)$ to $(m_{i},m_{j})$;
(n) for $i,j=h+1,\ldots,g$ there is an edge from $(i,j)$ to $(n_{j},n_{i})$.
We say that a vertex $(i,j)$ is marked if $(i,j)$ satisfies condition (3) or (4) of Lemma 4.5, that is, if $j\leqslant h<i$ or $(i,j)\in P\times Q$.
Lemma 4.6
The following properties hold:
(1) if there is an edge from $(i_{1},j_{1})$ to $(i_{2},j_{2})$, then $i_{1}-j_{1}$ and $i_{2}-j_{2}$ have the same sign (so the connected components ‘do not cross the diagonal’);
(2) every vertex has at most one incoming edge and at most one outgoing edge;
(3) the connected components of the graph $\mathcal{G}_{\varphi}$ are either cycles $(i_{1},j_{1})\rightarrow\ldots\rightarrow(i_{r},j_{r})\rightarrow(i_{1},j_{1})$ or chains $(i_{1},j_{1})\rightarrow\ldots\rightarrow(i_{s},j_{s})$ (without repetition of vertices);
(4) for every ‘diagonal’ vertex $(i,i)$ there is exactly one incoming and one outgoing edge; the edge starting at a diagonal vertex ends in a diagonal vertex;
(5) the connected components of the set of diagonal vertices are cycles;
(6) no edge originates at $(i,j)$ if and only if $i\leqslant h<j$ or $j\leqslant h<i$;
(7) there are exactly h(gh)h(g-h) chains above (respectively, below) the diagonal;
(8) no cycle contains a marked vertex;
(9) every chain below the diagonal ends with a marked vertex;
(10) every marked vertex above the diagonal is the start point of a chain;
(11) the number of chains without marked vertices is $\sum_{i=1}^{h}(m_{i}-i)$.
Proof. For a type (m) edge, $i<j$ implies $m_{i}<m_{j}$ because the function $i\mapsto m_{i}$ is order-preserving. For a type (n) edge, $i<j$ implies $n_{j}<n_{i}$ because the function $i\mapsto n_{i}$ is order-inversing. This proves (1). Properties (2), (4), (6) are immediate from the construction. Property (3) is a consequence of (2). Property (5) is a consequence of (4). A chain is uniquely determined by its end point $(i,j)$, which must be such that $i\leqslant h<j$ or $j\leqslant h<i$ by (6). Moreover, any such pair $(i,j)$ is the end point of some chain (which may have length 1). This implies (7). For (8), we note that a marked vertex is either $j\leqslant h<i$, in which case there is no outgoing edge, or $(i,j)\in P\times Q$, in which case there is no incoming edge. Property (9) is consequence of (6). Every marked vertex above the diagonal is of the form $(i,j)$, where $(i,j)\in P\times Q$, so there is no incoming edge, thus $(i,j)$ is the start point of a chain. This proves (10).
Let us prove (11). All marked points are either start or end points of chains, so a chain has no marked points if and only if its start point and its end point are not marked. This means that it starts at a vertex $(i,j)\in Q\times P$, and is above the diagonal. Thus a chain has no marked points if and only if it starts at a vertex $(i,j)\in Q\times P$ such that $j>i$. The number of such pairs $(i,j)$ is
$\sum_{i=1}^{h}|\{n\in Q|n<m_{i}\}|=\sum_{i=1}^{h}(m_{i}-i).$
This finishes the proof of the lemma. $\Box$
End of proof of Theorem 4.1. Let us prove ($\star$). By Lemma 4.5 we need to compute the space of $k$-valued functions $a_{ij}$ on the set of vertices of the directed graph $\mathcal{G}_{\varphi}$ such that if $(i,j)$ is a marked vertex then $a_{ij}=0$, and if there is an edge from $(i_{1},j_{1})$ to $(i_{2},j_{2})$ then $a_{i_{2},j_{2}}=a_{i_{1},j_{1}}^{p}$. This space is the direct sum of spaces given by the connected components of $\mathcal{G}_{\varphi}$.
The directed graph $\mathcal{G}_{\varphi}$ has the three kinds of connected components: cycles (which contain no marked vertices), chains with marked vertices, and chains with no marked vertices. The vector space associated to a cycle with $n$ vertices is isomorphic to ${\mathbb{F}}_{p^{n}}$. The vector space associated to a chain with marked vertices is 0. Finally, the vector space associated to a chain without marked vertices is isomorphic to $k$. The proof finishes by an application of Lemma 4.6 (11). $\square$
Corollary 4.7
Let $A$ be a principally polarised abelian variety over an algebraically closed field $k$ of characteristic $p\geqslant 3$. Let $r={\rm{dim}}({\bf U}_{A}[p])$. Then there exists an abelian variety $B$ over $k$ of dimension at most $r+1$ such that ${\bf U}_{A}\cong{\bf U}_{B}$. In particular, we have
${\rm{dim}}({\bf U}_{A})\leqslant r(r+1)/2\quad\text{and}\quad\operatorname{\mathnormal{p}-exp}({\bf U}_{A})\leqslant r.$
Proof. Let $f$ be the $p$-rank of $A$. Since $k$ is perfect, the $p$-divisible group $A[p^{\infty}]$ is isomorphic to the product of $p$-divisible groups of slopes 0 and 1, and the $p$-divisible group of an abelian variety $B$ of dimension gfg-f. By Corollaries 2.12 and 2.13 this implies that ${\bf U}_{A}\cong{\bf U}_{B}$.
By [O01, (9.7)], $f$ is the maximal value of $i$ such that $\varphi(i)=i$, i.e., the sequence $\varphi(i)$ starts with the sequence $1,2,\ldots,f-1,f,f$. Thus $f$ is the maximal value of $i$ such that $m_{i}=i$. In particular, we have $m_{i}-i\geqslant 1$ for $i\geqslant f+1$. Now Theorem 4.1 gives
$r=\frac{a(a-1)}{2}+\sum_{i=1}^{g-a}(m_{i}-i)\geqslant\frac{a(a-1)}{2}+g-a-f=\frac{a(a-3)}{2}+g-f\geqslant g-f-1,$
where we used that $a(a-3)/2\geqslant-1$ for all $a\in{\mathbb{Z}}$. Now (2) gives ${\rm{dim}}({\bf U}_{A})={\rm{dim}}({\bf U}_{B})\leqslant(g-f)(g-f-1)/2\leqslant r(r+1)/2$. $\Box$
Example 4.8
It is instructive to explore what Corollary 4.7 gives for small values of $r$. Suppose that ${\bf U}_{A}$ is isogenous to ${\bf W}_{n_{1}}\times\ldots\times{\bf W}_{n_{r}}$, where $n_{1}\leqslant\ldots\leqslant n_{r}$.
(1) If $r=1$, then $n_{1}=1$.
(2) If $r=2$, then $(n_{1},n_{2})$ is either $(1,1)$ or $(1,2)$.
(3) If $r=3$, then $n_{1}+n_{2}+n_{3}\leqslant 6$. There are two possible cases: ${\rm{dim}}(B)=3$ and ${\rm{dim}}(B)=4$. In the first case we have $(n_{1},n_{2},n_{3})=(1,1,1)$. In the second case, we have $n_{i}\leqslant 3$ by Theorem 3.7. Listing all possibilities for the Newton polygon of $B$, we see that $(n_{1},n_{2},n_{3})=(1,1,2)$ is the only possibility in the non-supersingular case besides $(1,1,1)$. In the supersingular case we have $n_{1}+n_{2}+n_{3}=6$. In view of Proposition 3.11 we expect that the case $(n_{1},n_{2},n_{3})=(1,2,3)$ does occur, but we are not sure if $(n_{1},n_{2},n_{3})=(2,2,2)$ does.
Corollary 4.9
The following table gives all possible isogeny classes of ${\bf U}_{A}$ for principally polarised abelian threefolds over an algebraically closed field $k$ of characteristic $p\geqslant 3$ of a given Ekedahl–Oort type:
E-O type $\varphi$ $a$ $p$-rank Newton polygon ${\bf U}_{A}$ up to isogeny
$0,0,0$ 3 0 supersingular $({\bf G}_{a,k})^{3}$
$0,0,1$ 2 0 supersingular $({\bf G}_{a,k})^{3}$
$0,1,2$ 1 0 supersingular ${\bf G}_{a,k}\times{\bf W}_{2}$
$0,1,2$ 1 0 $1/3$-type $({\bf G}_{a,k})^{2}$
$0,1,1$ 2 0 $1/3$-type $({\bf G}_{a,k})^{2}$
$1,1,1$ 2 1 almost supersingular ${\bf G}_{a,k}$
$1,1,2$ 1 1 almost supersingular ${\bf G}_{a,k}$
$1,2,2$ 1 2 almost ordinary 0
$1,2,3$ 0 3 ordinary 0
Proof. The Ekedahl–Oort type of ${\bf U}_{A}[p]$ determines $a$, the $p$-rank, and ${\rm{dim}}({\bf U}_{A}[p])$, whereas the Newton polygon determines ${\rm{dim}}({\bf U}_{A})$.
Ekedahl–Oort types $(0,0,0)$ and $(1,2,3)$ were dealt with in Example 4.2.
If $a=2$, then $\varphi$ can be $(0,0,1)$, $(0,1,1)$, and $(1,1,1)$. A result of Oort [O01, Theorem (8.3)] says that in the first case $A$ is supersingular, in the second case $A$ of the $1/3$-type, and in the third case $A$ is almost supersingular.
If $a=1$, then $\varphi$ can be $(0,1,2)$, $(1,1,2)$, and $(1,2,2)$. We know from [O01, (9.7)] that the $p$-rank $f$ of $A$ is the maximal value of $i$ such that $\varphi(i)=i$. This implies that in the last case, the $p$-rank of $A$ is 2. Thus we have an isomorphism of group $k$-schemes $A[p]\cong({\mathbb{Z}}/p)^{2}\times(\mu_{p})^{2}\times G$, where $G$ is local-local. It follows that $A$ is almost ordinary. The same argument gives that in the second case the $p$-rank of $A$ is 1. Then $A[p]\cong{\mathbb{Z}}/p\times\mu_{p}\times G$, where $G$ is local-local, and it follows that $A$ is almost supersingular. Finally, if $\varphi$ is $(0,1,2)$, then the $p$-rank of $A$ is zero, so the Newton polygon of $A$ is either supersingular or of $1/3$-type. $\Box$
5 Open questions
The main question of this paper is
Question 5.1
Let $A$ be an abelian variety of dimension $g$ over an algebraically closed field of characteristic $p>0$. What are the possible integers $n_{1}\leqslant\ldots\leqslant n_{r}$ such that ${\bf U}_{A}$ is isogenous to $\prod_{i=1}^{r}{\bf W}_{n_{i}}$?
We can ask more specific questions about various particular types of abelian varieties.
The reader will notice that our proof of Theorem 3.7 in the non-supersingular case bounded the dimension of ${\rm{End}}_{\mathbb{E}}(M^{\vee}/p^{m},M/p^{m})^{\rm skew}$ by the dimension of the larger quasi-algebraic group ${\rm{End}}_{\mathbb{E}}(M^{\vee}/p^{m},M/p^{m})$. It is thus likely that in this case the bound g1g-1 for the $p$-exponent of ${\bf U}_{A}$ is not the best possible.
Question 5.2
Let $A$ be a non-supersingular abelian variety of dimension $g\geqslant 2$ over an algebraically closed field of characteristic $p>0$. Is it true that $p^{g-2}{\bf U}_{A}=0$?
Assuming $p\neq 2$, this holds for $g\leqslant 3$ (Proposition 3.15) and also for $g=4$ in the principally polarised case, as one sees by arguing similarly to Example 4.8.
In the proof of Proposition 3.11 we used a supersingular abelian variety $A$ with Dieudonné module $M={\mathbb{E}}/(F^{g}-V^{g}){\mathbb{E}}$. We note that the $a$-number of $A$ equals ${\rm{dim}}_{k}(M/(FM+VM))=1$, that is, $A$ is supergeneral.
Question 5.3
Let $A$ be a supersingular abelian variety of dimension $g\geqslant 3$ over an algebraically closed field of characteristic $p>0$ such that the $a$-number of $A$ is at least $2$ (that is, $A$ is not supergeneral). Is it true that $p^{g-2}{\bf U}_{A}=0$?
Here we must exclude the case $g=2$ because for a supersingular elliptic curve $E$ we have $a(E^{2})=2$ whereas ${\bf U}_{E^{2}}\cong{\bf G}_{a,k}$.
Finally, one can enquire whether the description of the isogeny class of ${\bf U}_{A}$ in Proposition 3.11 remains true for all supergeneral abelian varieties (this is [YY25, Conjecture 5.2.1]). In general it would be interesting to understand which numerical invariants of $M$ completely determine the isogeny type of ${\bf U}_{A}$.
Question 5.4
Consider the abelian varieties $A$ over an algebraically closed field $k$ of characteristic $p>0$ with a fixed Newton polygon and $a$-number 11. Is the isogeny type of ${\bf U}_{A}$ independent of $A$?
For $g\leqslant 3$ the answers to Questions 5.3 and 5.4 are positive, see Proposition 3.15. The positive answer to Question 5.4 for supergeneral varieties would be in agreement with Proposition 3.13, which says that in that case ${\rm{dim}}({\bf U}_{A}[p])=g-1$.
In the appendix below, we list the isogeny types of ${\bf U}_{A}$ for some abelian varieties $A$ with $a$-number 11. The following curious question seems to have a positive answer for all the examples at our disposal.
Question 5.5
Does the isogeny type $n_{1}\leqslant\ldots\leqslant n_{r}$ of ${\bf U}_{A}$ always contain a subsequence $1,2,\ldots,n_{r}-1,n_{r}$?
Appendix A Numerical computations of isogeny types
Below is a list of isogeny types of ${\bf U}_{A}$ for a selected set of abelian varieties $A$.
$g$ Newton polygon $\nu$ ${\rm{dim}}({\bf U}_{A_{\nu}})$ Isogeny type of ${\bf U}_{A_{\nu}}$
3 $\left[\frac{1}{2},\frac{1}{2},\frac{1}{2}\right]$ 3 $\left[1,2\right]$
$\left[\frac{1}{3},\frac{2}{3}\right]$ 2 $\left[1,1\right]$
4 $\left[\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2}\right]$ 6 $\left[1,2,3\right]$
$\left[\frac{1}{3},\frac{1}{2},\frac{2}{3}\right]$ 4 $\left[1,1,2\right]$
$\left[\frac{1}{4},\frac{3}{4}\right]$ 3 $\left[1,1,1\right]$
5 $\left[\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2}\right]$ 10 $\left[1,2,3,4\right]$
$\left[\frac{2}{5},\frac{3}{5}\right]$ 8 $\left[1,2,2,3\right]$
$\left[\frac{1}{3},\frac{1}{2},\frac{1}{2},\frac{2}{3}\right]$ 7 $\left[1,1,2,3\right]$
$\left[\frac{1}{4},\frac{1}{2},\frac{3}{4}\right]$ 5 $\left[1,1,1,2\right]$
$\left[\frac{1}{5},\frac{4}{5}\right]$ 4 $\left[1,1,1,1\right]$
6 $\left[\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2}\right]$ 15 $\left[1,2,3,4,5\right]$
$\left[\frac{2}{5},\frac{1}{2},\frac{3}{5}\right]$ 12 $\left[1,2,2,3,4\right]$
$\left[\frac{1}{3},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{2}{3}\right]$ 11 $\left[1,1,2,3,4\right]$
$\left[\frac{1}{3},\frac{1}{3},\frac{2}{3},\frac{2}{3}\right]$ 10 $\left[1,2,2,2,3\right]$
$\left[\frac{1}{4},\frac{1}{2},\frac{1}{2},\frac{3}{4}\right]$ 8 $\left[1,1,1,2,3\right]$
$\left[\frac{1}{5},\frac{1}{2},\frac{4}{5}\right]$ 6 $\left[1,1,1,1,2\right]$
$\left[\frac{1}{6},\frac{5}{6}\right]$ 5 $\left[1,1,1,1,1\right]$
Here we consider abelian varieties of dimension $g$ with $p$-rank 0 and $a$-number 11. Their Dieudonné module is then isomorphic to
${\mathbb{E}}/(a_{0}F^{g}+a_{1}F^{g-1}+\dots+a_{2g}V^{g}){\mathbb{E}},$ (27)
for some $a_{i}\in W$, where $a_{0},a_{2g}\in W^{\times}$.
Let $\nu$ be a symmetric Newton polygon which has no slopes 0 or $1$. If $M_{\nu}$ is the Dieudonné module of the form (27) where
[…]
then $M_{\nu}$ has Newton polygon $\nu$ by [LNV13, Lemma 5.8]. We take $A_{\nu}$ to be any abelian variety with Dieudonné module $M_{\nu}$.
The table contains the isogeny types of ${\bf U}_{A_{\nu}}$ for $g\leq 6$, when $p=3,5,7$ (the result does not depend on such choice of $p$). The computation was done in SageMath v10.7 using the implementation of Witt vectors in [Mu25]. The code can be found at https://github.com/liviabasilico/brauer-p-torsion. The principle of the SageMath script is to find the equations that define ${\rm{Hom}}_{\mathbb{E}}(M^{\vee}/p^{n},M/p^{n})^{\rm skew}$ as a subvariety of ${\rm{Hom}}_{W_{n}}(M^{\vee}/p^{n},M/p^{n})$ and hence compute its dimension.